### v1

# The Class 0 protostars in Orion: Characterizing the properties of their magnetized envelopes

Abstract
We present a study connecting the physical properties of protostellar envelopes to the morphology of the envelope-scale magnetic field. We use the ALMA polarization observations of 55 young prtostars at 0.87 mm on $\sim 400-3000$ au scales from the B-field Orion Protostellar Survey (BOPS) to infer the envelope-scale magnetic field and both dust and gas emission on comparable scales to measure the envelope properties. We find that the protostellar envelopes with compact polarized dust emission tend to have lower envelope masses, than the sources with more extended envelopes. We also find that protostars showing hourglass-field morphologies tend to have lower velocity dispersions in their envelopes, whereas systems with spiral-field morphologies have higher velocity dispersion. Combining with the disk properties taken from the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM) survey, we connect envelope properties to fragmentation. Our results suggest that envelope mass may not correlate with fragmentation, whereas turbulence appears to promote fragmentation. On the other hand, we find that fragmentation is suppressed in systems with pinched magnetic fields, suggesting that the magnetic field play a role on providing additional support against gravitational collapse, and the formation of an hourglass-like field may coincide with enhanced magnetic braking that removes angular momentum and hinders the formation of embedded disks. Nevertheless, significant misalignment between magnetic field and outflow axes tends to reduce magnetic braking, leading to the formation of larger disks.
Star formation - star forming regions - magnetic fields - interstellar magnetic fields - circumstellar envelopes - circumstellar disks
1 Introduction
Fragmentation is thought to be a critical process in the formation of star and stellar systems (e.g., Matsumoto et al., 2003). Initially, turbulence induces chaotic motions and density fluctuations across various scales, compressing the gas in certain regions and leading to localized density enhancements (e.g., Mac Low et al., 2004). Within these dense regions, gravitational forces overcome thermal pressure if the mass exceeds the Jeans mass. As the cloud collapses, large-structure fragment into smaller, dense cores, each of which form individual stars or multiple star systems (e.g., McKee et al., 2007). This hierarchical fragmentation continues until individual cores reach densities high enough to resist further fragmentation (e.g., Beuther et al., 2015). However, the specific mechanisms by which a cloud core fragments to ultimately form a multiple system or small cluster remain uncertain.
Theoretical and numerical studies (e.g., Maury et al., 2022; Offner et al., 2023) suggest that cores with some angular momentum are prone to fragmentation into multiple objects, indicating the need for additional mechanisms to reduce the fragmentation to the observed levels. Since non-ideal MHD effects do not inhibit protostellar multiplicity (e.g., Wurster et al., 2019), further research points to turbulence and magnetic fields (hereafter, B-fields) as playing a pivotal role in the fragmentation process (e.g., Padoan et al., 2001; Price et al., 2009). For instance, compressive turbulence promotes fragmentation and the formation of structures with high density contrasts, whereas solenoidal turbulence (e.g., Federrath et al., 2010; Girichidis et al., 2011) suppresses fragmentation and maintains a smoother density structure within the cloud. On the other hand, the ratio of rotational to magnetic energy has been found to influence core fragmentation (e.g., Machida et al., 2008). Moreover, B-field tends to suppress fragmentation by providing additional support against gravitational collapse, and/or through magnetic braking, which carries away the angular momentum from the inner regions of the core (e.g., Commerçon et al., 2011; Guszejnov et al., 2021; Mignon-Risse et al., 2021).
Dust polarization is the most commonly used tracer of B-fields in star-forming regions, as “radiative torques” (e.g., Hoang et al., 2009; Andersson et al., 2015) tend to align spinning and elongated dust grains with their long axes perpendicular to the ambient field direction. Over the last decades, dust polarization observations carried out with (sub)millimeter interferometers have increasingly proven effective in mapping B-fields at the scales of cores ($\sim 10^{4}$ au) and envelopes ($\sim 10^{2}$ to $10^{4}$ au) (e.g. Girart et al., 1999, 2013; Cox et al., 2018; Galametz et al., 2018; Hull et al., 2019; Le Gouellec et al., 2020; Cortés et al., 2021). Theory predicts that B-field regulates the collapse of clouds, while ambipolar diffusion allows the formation of “supercritical” dense cores, where the gravitational force dominates over the magnetic support. As cloud cores collapse, field lines are dragged inward to be drawn into pinched morphology. This has been observed in various studies (e.g., Girart et al., 2006, 2009; Qiu et al., 2014; Beltrán et al., 2019; Le Gouellec et al., 2019; Hull et al., 2020). In the case of flux freezing, the pinched geometry in the pseudo-disk strengthens the B-field and carries away angular momentum, leading to catastrophic magnetic braking and potentially preventing the disk formation (e.g., Galli et al., 2006). However, disk may form if the magnetic and rotational axes are misaligned (e.g., Joos et al., 2012; Li et al., 2013; Machida et al., 2020). To statistically test these scenarios related to fragmentation and star/disk formation, it is crucial to understand the connection between envelope properties and the resulting fragmentation.
Recently, the first results from the B-fields Orion protostellar survey (BOPS, with a resolution of $\sim 1^{\prime\prime}$, corresponding to $\sim$ 400 au, Huang et al., 2024) presented the largest polarization study of B-fields in protostellar envelopes using dust polarization observations of 61 protostars in Orion. We successfully mapped the B-fields morphology at scales ranging from several hundred to a few thousand au, and classified the protostars into several different morphological categories: standard hourglass (with the main axis coincident with the outflow axis), rotated hourglass (similar to the standard hourglass but rotated by $\sim 90^{\circ}$, with its axis perpendicular to the outflow), spiral configuration, and complex B-field structure. In addition, observations from a major Orion interferometric study, known as the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM, Tobin et al., 2020), have identified protostellar disks in the Orion molecular clouds with a resolution of $\sim 0\farcs 1$ (corresponding to $\sim$ 40 au). This study has provided data on the radius and mass of these disks. Since all BOPS protostars were originally observed at high resolution in the VANDAM survey, the combination of these datasets enables a comprehensive test of the magnetic braking model and allows for a more effective linkage of magnetic and turbulent properties to the characteristics of protostellar disks and multiplicity.
This paper builds upon the first results of BOPS (Huang et al., 2024) and is organized as follows: Section 2 introduces the observations and data reduction processes. Section 3 presents the methods for calculation of physical parameters, the velocity dispersion, the angle dispersion of the B-field, and the fragmentation level. A detailed discussion of the correlation between these parameters is provided in Section 4. Finally, the conclusions are summarized in Section 5.
2 Observations
The 870 $\mu$m (Band 7) dust polarization observations were taken with ALMA in 2019 November, and the spatial resolution and the Largest Angular Scale (LAS) are $\sim$1″and $\sim$8″, corresponding to a linear resolution of $\sim$400 and $\sim$3200 au at a distance of $\sim$400 pc (see Huang et al. 2024 for further details on the observation setup).
[…]
For the C17O (3–2) line data, we used the CASA uvcontsub task to perform continuum fitting and subtraction in the $u$-$v$ plane. We then utilized the tclean task to produce Stokes I image cubes using self-calibrated, continuum-subtracted data for each spectral window. The C17O (3–2) line, optically thin in dense regions (with spectral resolution of $\sim$ 0.87 km s-1, Huang et al., 2024), was employed to trace the velocity field of the envelope.
3 Analysis
Of the original 61 protostars presented in Huang et al. (2024), we excluded 6 objects due to a lack of data on either their polarization (HOPS-87S and ONC1N-8-S) or the embedded star/disk (e.g., insufficient information to determine their bolometric luminosity and embedded disk properties, ONC1N-4-5-EN, ONC1N-4-5-ES, ONC1N-6-7, and ONC1N-8-N). Additionally, 16 objects exhibited polarization measurements consistent with self-scattering in their disks and no B-field detections in the envelopes. These 16 objects had uniform polarization morphologies that aligned with the disk minor axes, a tell-tale sign of self-scattering (Kataoka et al., 2015, 2017; Stephens et al., 2017). Furthermore, 7 protostellar envelopes exhibited unresolved polarization measurements. This leaves us with a sample of 32 protostars with robust B-field measurements in their envelopes and complementary star and disk information. Among these remaining protostars, 6 exhibit standard hourglass field structures, 11 show rotated hourglass field morphologies, 9 have spiral magnetic fields, and 6 have complex or disordered field morphologies. The following sections presents the estimation of physical envelope parameters, velocity dispersion, the angle dispersion of B-field, and fragmentation level. These properties will then be compared to the disk properties to understand which factors could potentially influence the envelope fragmentation and whether the initial misalignment between the B-field and outflows could result in reduced magnetic braking during disk formation.
3.1 Physical Envelope Properties
[…]
$M_{\rm dust}=\frac{S_{\nu}d^{2}}{\kappa_{\nu}B_{\nu}(T_{\rm dust})}$ (1)
where $d$ is the distance to the protostar as obtained from Huang et al. (2024), $\nu$ is the frequency corresponding to the observed wavelength, $B_{\nu}(T_{\rm dust})$ is the Planck function at the dust temperature of $T_{\rm dust}$, and $\kappa_{\nu}$ is the dust opacity. We adopt $\kappa_{870\mu{\rm m}}\approx 1.84\ {\rm cm}^{2}\ {\rm g}^{-1}$, which is appropriate for dusty particles with ice mantles at densities of approximately $10^{6}$ cm-3, as reported by Ossenkopf et al. (1994).
Table 1: Envelope and disk parameters for the BOPS sample\$$\dagger$\$$\dagger$\$footnotemark:
Name $L_{\rm{bol}}$ $S_{\nu}$ $\overline{T}_{\rm env}$ $T_{\rm c}$ $M_{\rm env}$ $N_{\rm gas}$ $n_{\rm gas}$ $\sigma_{\rm nth}$ $\delta\phi$ $N_{F}$ $M_{\rm disk}$bbfootnotemark: $R_{\rm disk}$bbfootnotemark: $\Delta\theta$ccfootnotemark:
[…]
Standard
HOPS-11 9.00 0.40 26.1 16.7 0.26 1.3E+23 5.4E+06 0.72 / 1.00 135 41.9 14.0
HOPS-87aafootnotemark: 36.49 3.89 37.0 17.6 1.46 7.3E+23 3.0E+07 0.70 17.6$\pm$1.4 1.00 628 124.1 49.1
HOPS-359 10.00 0.91 26.8 16.1 0.62 3.1E+23 1.3E+07 0.65 29.2$\pm$2.3 1.00 90 62.0 15.6
HOPS-395 0.50 0.58 20.2 11.7 0.49 2.4E+23 1.0E+07 0.70 15.0$\pm$2.0 2.00 24 14.2 0.1
HOPS-400 2.94 1.18 21.5 13.7 1.01 5.0E+23 2.1E+07 0.82 18.2$\pm$1.9 2.00 107 48.7 14.3
HOPS-407 0.71 0.61 20.2 13.7 0.59 2.9E+23 1.2E+07 0.64 10.5$\pm$1.0 1.00 83 64.1 16.5
Rotated
HH270IRS 7.70 0.49 25.1 16.6 0.35 1.7E+23 7.2E+06 0.77 33.8$\pm$4.1 2.00 161 92.3 68.3
HOPS-78 8.93 0.81 26.0 19.8 0.48 2.4E+23 9.9E+06 0.77 40.2$\pm$4.3 2.25 188 147.1 86.0
HOPS-168 48.07 0.51 39.6 18.1 0.17 8.4E+22 3.5E+06 0.90 28.4$\pm$3.9 2.00 65 65.8 63.4
HOPS-169 3.91 0.67 22.2 15.7 0.47 2.3E+23 9.7E+06 0.63 32.9$\pm$3.9 1.00 119 53.8 51.3
HOPS-288 135.47 1.60 51.4 18.8 0.43 2.1E+23 8.9E+06 1.38 24.5$\pm$2.9 3.00 263 144.3 80.0
HOPS-310 13.83 0.58 29.1 17.7 0.33 1.6E+23 6.8E+06 0.95 / 1.00 263 101.3 89.9
HOPS-317Saafootnotemark: 4.90 2.33 22.9 14.2 1.95 9.7E+23 4.0E+07 0.64 / 1.00 1059 269.1 83.4
HOPS-325 6.20 0.28 23.9 22.5 0.22 1.1E+23 4.6E+06 0.65 / 1.00 59 34.9 51.0
HOPS-370 360.86 1.08 65.6 25.4 0.21 1.0E+23 4.3E+06 1.22 33.8$\pm$3.1 1.00 114 112.8 86.8
HOPS-401 0.61 0.37 20.2 14.2 0.37 1.9E+23 7.7E+06 0.63 / 1.00 186 74.7 78.8
HOPS-409 8.18 0.38 25.4 22.5 0.23 1.2E+23 4.8E+06 0.78 17.0$\pm$2.5 1.00 111 112.1 73.3
Spiral
HOPS-96 6.19 0.44 23.8 19.9 0.29 1.4E+23 6.0E+06 0.57 / 1.00 135 41.4 52.4
HOPS-124 58.29 1.43 41.6 17.8 0.48 2.4E+23 9.9E+06 1.51 / 2.00 438 119.3 80.5
HOPS-182 71.12 1.43 43.7 19.7 0.42 2.1E+23 8.7E+06 1.15 44.8$\pm$4.5 2.25 238 69.7 32.7
HOPS-303 1.49 0.42 20.6 17.3 0.37 1.9E+23 7.8E+06 0.73 / 1.00 207 82.9 53.9
HOPS-361N 478.99 1.70 70.4 23.3 0.36 1.8E+23 7.5E+06 0.92 41.3$\pm$3.2 2.75 202 173.6 45.1
HOPS-361S 478.99 1.10 70.4 23.3 0.23 1.2E+23 4.8E+06 0.88 43.1$\pm$3.8 4.00 139 96.1 40.2
HOPS-384 1477.95 1.98 93.3 33.9 0.28 1.4E+23 5.8E+06 1.02 30.8$\pm$2.2 2.50 32 36.5 67.9
HOPS-403aafootnotemark: 4.14 0.89 22.3 14.8 0.77 3.8E+23 1.6E+07 1.12 / 2.00 479 140.4 42.3
HOPS-404 0.95 0.53 20.4 13.3 0.53 2.6E+23 1.1E+07 0.70 / 1.00 389 59.9 32.0
Complex
HOPS-12W 7.31 0.48 24.8 16.5 0.30 1.5E+23 6.1E+06 0.63 / 2.00 20 20.5 75.3
HOPS-88 15.81 0.70 30.0 18.2 0.34 1.7E+23 7.1E+06 0.76 33.5$\pm$4.8 1.00 71 34.0 53.2
HOPS-373E 5.32 0.53 23.2 15.4 0.43 2.2E+23 9.0E+06 0.68 / 1.50 79 69.8 /
HOPS-373W 5.32 0.64 23.2 15.4 0.52 2.6E+23 1.1E+07 0.71 / 1.50 70 36.0 /
HOPS-398 1.01 0.62 20.4 17.2 0.55 2.8E+23 1.2E+07 0.47 / 1.00 272 76.9 /
HOPS-399 6.34 1.79 24.0 17.2 1.28 6.3E+23 2.6E+07 0.70 46.4$\pm$4.2 1.00 197 66.2 67.6
Unresolved
HOPS-164 0.58 0.24 20.2 13.4 0.19 9.7E+22 4.0E+06 0.77 / 1.00 134 70.6 /
HOPS-224 2.99 0.36 21.5 13.1 0.34 1.7E+23 7.1E+06 0.78 / 1.00 237 107.7 /
HOPS-250 6.79 0.20 24.3 13.5 0.16 7.8E+22 3.3E+06 0.92 / 1.00 134 117.2 /
HOPS-317N 4.76 0.11 22.9 14.2 0.09 4.5E+22 1.9E+06 0.67 / 1.00 52 34.5 /
HOPS-341 2.07 0.24 21.0 14.7 0.23 1.1E+23 4.8E+06 0.63 / 1.00 28 54.5 /
HOPS-402 0.55 0.51 20.2 13.3 0.51 2.5E+23 1.1E+07 0.66 / 1.00 622 156.6 /
HOPS-408 0.52 0.31 20.2 12.2 0.27 1.3E+23 5.6E+06 0.94 / 1.00 125 93.5 /
Scattering
HH212M 14.00 0.52 29.1 18.2 0.32 1.6E+23 6.6E+06 0.73 / 1.00 103 78.6 /
HOPS-10 3.33 0.17 21.8 16.5 0.12 6.2E+22 2.6E+06 0.68 / 1.00 53 46.8 /
HOPS-12E 7.31 0.33 24.8 16.5 0.21 1.0E+23 4.3E+06 0.78 / 2.50 62 58.7 /
HOPS-50 4.20 0.32 22.4 18.8 0.23 1.2E+23 4.8E+06 1.01 / 1.00 173 82.8 /
HOPS-53 26.42 0.37 34.1 21.2 0.15 7.6E+22 3.2E+06 0.74 / 1.00 33 21.6 /
HOPS-60 21.93 0.55 32.6 22.9 0.24 1.2E+23 5.0E+06 0.87 / 1.00 153 62.7 /
HOPS-81 1.24 0.14 20.5 17.9 0.11 5.6E+22 2.4E+06 0.74 / 1.00 66 32.4 /
HOPS-84 49.11 0.59 39.8 20.5 0.20 1.1E+23 4.2E+06 0.92 / 2.00 159 87.4 /
HOPS-153 4.43 0.22 22.5 14.0 0.16 7.8E+22 3.3E+06 0.92 / 1.00 103 129.1 /
HOPS-203N 20.44 0.42 32.0 15.7 0.18 9.0E+22 3.8E+06 0.86 / 3.00 31 58.6 /
HOPS-203S 20.44 0.38 32.0 15.7 0.16 8.1E+22 3.4E+06 0.71 / 3.00 33 30.9 /
HOPS-247 3.09 0.37 21.6 13.1 0.34 1.7E+23 7.1E+06 0.84 / 1.00 227 85.3 /
HOPS-340 1.85 0.21 20.8 14.7 0.20 1.0E+23 4.2E+06 0.70 / 1.00 60 41.7 /
HOPS-354 6.57 0.34 24.2 / 0.18 9.1E+22 3.8E+06 0.90 / 1.00 30 24.4 /
HOPS-358 24.96 0.44 33.7 17.6 0.22 1.1E+23 4.6E+06 1.23 / 1.00 111 136.3 /
HOPS-383 7.83 0.22 25.2 20.3 0.14 6.9E+22 2.9E+06 0.89 / 1.00 96 61.7 /
Column 2 to 14 present the bolometric luminosity ($L_{\rm{bol}}$), total continuum flux density at 345 GHz ($S_{\nu}$), averaged envelope dust temperature ($\overline{T}_{\rm env}$), cloud dust temperature ($T_{\rm c}$), envelope gas mass ($M_{\rm env}$), gas column density ($N_{\rm gas}$), gas number density ($n_{\rm gas}$), non-thermal velocity dispersion ($\sigma_{\rm nth}$), B-field angle dispersion ($\delta\phi$), fragmentation level ($N_{F}$), disk mass ($M_{\rm disk}$), disk radius ($R_{\rm disk}$), and the misalignment between outflow and B-field ($\Delta\theta=|\theta_{O}-\theta_{B}|$), respectively. “Standard”, “Rotated”, “Spiral”, “Complex” indicate protostars with different types of B-field structure, “Unresolved” means the B-field pattern is not sufficient to identify its morphology, while “Scattering” represents protostars with compact polarized emission, which is thought to be produced by self-scattering (Huang et al., 2024).$\dagger$\$$\dagger$\$footnotemark: Typical uncertainties for the different parameters are: 10% for $S_{\nu}$; 20–30% for $\overline{T}_{\rm env}$; 30–40% for $M_{\rm env}$, $N_{\rm gas}$, and $n_{\rm gas}$; 0.37 km s-1 for $\sigma_{\rm nth}$.aafootnotemark: The disk emission is contaminated by the envelope dust emission.bbfootnotemark: The disk mass, radius, and their uncertainties are obtained from Tobin et al. (2020).ccfootnotemark: The misalignment between outflow and B-field and its uncertainty are obtained from Huang et al. (2024).
Table 1: Envelope and disk parameters for the BOPS sample\$$\dagger$\$$\dagger$\$footnotemark:
In the case of a dusty cloud heated by internal stellar radiation, the radial temperature profile can be expressed as a power law: $T_{\rm dust}\propto r^{-2/(4+\beta)}$ (Kenyon et al., 1993), where $\beta$ is the dust opacity spectral index. Girart et al. (2009) adopted a value of $\beta=1.5$ when studying low-mass protostars at 1000 au scales. While Kwon et al. (2009) suggested that $\beta$ should be around or smaller than 1 for Class 0 stage. Tobin et al. (2020) fitted a temperature power-law index of -0.46 for hundreds-au-size disks within the Orion star-forming protostellar system, corresponding to $\beta\sim 0.35$, scaled by luminosity. However, this profile shows a slight decline in the outer 100 au regions due to backheating from the surrounding envelope (Tobin et al., 2020), it is subsequently recommended to refine the index to -0.50 (Sheehan et al., 2022). For this study, we use the temperature model proposed by Tobin et al. (2020) for Orion star-forming regions, with $\beta$ set to 1 (corresponding to the index of $\sim-0.40$) for the Class 0 protostellar envelope with a radius of 1200 au. Therefore, the dust temperature $T_{\rm dust}$ can be expressed as:
$T_{\rm dust}=T_{0}\biggl{(}\frac{L_{\rm bol}}{L_{\odot}}\biggr{)}^{0.25}\biggl{(}\frac{r}{50~{}{\rm au}}\biggr{)}^{-0.40}$ (2)
where $T_{0}=43$ K is the average temperature for a $\sim$ 1 $L_{\odot}$ protostar at a radius of $\sim$50 au (Whitney et al., 2003; Tobin et al., 2013, 2020), $r$ is the distance from the protostar, and $L_{\rm bol}$ is the bolometric luminosity taken from Tobin et al. (2020) and Federman et al. (2023). Given the clear detection of C17O emission in our sample, it is inferred that the dust temperature cannot fall below 20 K, as CO would otherwise be frozen onto the dust grains, making it undetectable (e.g., Jørgensen et al., 2015). The Equation 2 may yield very low dust temperatures (below 20 K) at the edge of the selected regions, especially when the bolometric luminosity $L_{\rm bol}$ is very low. In such cases, we use 20 K as the minimum dust temperature for pixels where the calculated temperature falls below this value. Excluding the extremely high temperature in the central pixel, we calculate the average envelope dust temperature $\overline{T}_{\rm env}$ within inner regions with a radius of $R=1200$ au.
Lombardi et al. (2014) used Planck and Herschel observations of Orion to construct and fit SEDs, and further to obtain a map of dust temperature. As shown in Figure 1, comparing the cloud dust temperature $T_{\rm c}$ (Lombardi et al., 2014) with the envelope dust temperature $\overline{T}_{\rm env}$ (this work, derived using our method described above), we find that 3 protostars have values of $\overline{T}_{\rm env}$ comparable to the dust temperature of the surrounding molecular cloud (near the grey dashed line in Figure 1). Among these, two protostars (HOPS-325 and HOPS-409) exhibit a rotated-hourglass structure of the B-field, while one (HOPS-81) shows compact polarized emission (self-scattering). The remaining protostars have $\overline{T}_{\rm env}$ values higher than $T_{\rm c}$ (above the grey dashed line in Figure 1). Given that the scales of molecular clouds in question are on the parsec level, which is much larger than the envelope scales we focus on, it is expected that $\overline{T}_{\rm env}$ should be comparable to or higher than $T_{\rm c}$, depending on whether the bolometric luminosity is low or high. One should note that $T_{\rm c}$ is not available for HOPS-354, as it is outside the region covered by Lombardi et al. (2014). However, the comparison between $\overline{T}_{\rm env}$ and $T_{\rm c}$ (see the right panel of Figure 1) indicates that $\overline{T}_{\rm env}$ is within a reasonable range for all protostars. Therefore, we use $\overline{T}_{\rm env}$ to estimate the envelope physical properties for HOPS-354.
Figure 1: The difference between derived dust temperature within envelope $\overline{T}_{\rm env}$ and cloud dust temperature $T_{\rm c}$ obtained from Lombardi et al. (2014). Grey dashed line indicates when $\overline{T}_{\rm env}=T_{\rm c}$.
The gas column density $N_{\rm gas}$ is estimated using the relation:
$N_{\rm gas}=\frac{M_{\rm env}}{\pi R^{2}\mu_{\rm gas}m_{\rm H}}$ (3)
where $R=1200$ au is the radius, $M_{\rm env}=\eta M_{\rm dust}$ is the envelope gas mass, $\eta=100$ is the gas-to-dust mass ratio, $\mu_{\rm gas}=2.37$ is the mean molecular weight per free particle (Kauffmann et al., 2008), and $m_{\rm H}$ is the mass of the hydrogen atom. For a spherical region, the average gas number density can be calculated as
$n_{\rm gas}=\frac{3N_{\rm gas}}{4R}$ (4)
Table LABEL:Tab:parameters summarizes the derived parameters for each protostar, including information from the literature. Column 2 to 8 of Table LABEL:Tab:parameters provide the following data: bolometric luminosity $L_{\rm bol}$, total flux density $S_{\nu}$, mean envelope dust temperature $\overline{T}_{\rm env}$, cloud dust temperature $T_{\rm c}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, and gas number density $n_{\rm gas}$. Figure 2(A–F) illustrates the distributions of these parameters. The median and mean values of the sample are as follows: $L_{\rm bol}$: 6.57 and 63.87 L, $S_{\nu}$: 0.51 and 0.71 Jy, $\overline{T}_{\rm env}$: 24.2 and 29.9 K, $M_{\rm env}$: 0.30 and 0.40 M, $N_{\rm gas}$: 1.5×$\times 10^{23}$ and 2.0×$\times 10^{23}$ cm-2, and $n_{\rm gas}$: 6.1×$\times 10^{6}$ and 8.2×$\times 10^{6}$ cm-3.
Figure 2: The stacked histogram of physical envelope parameters for protostars with different B-field structure. Panel A to H: bolometric luminosity $L_{\rm bol}$, mean dust envelope temperature $\overline{T}_{\rm d}$, total flux density $S_{\nu}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, gas number density $n_{\rm gas}$, non-thermal velocity dispersion $\sigma_{\rm nth}$ and angle dispersion of B-field $\delta\phi$, respectively. Dashed and solid lines indicate the mean and median values of the distribution.
3.2 Line Width
Assuming that non-thermal motions on envelope scales arise from turbulence, we use the line width of the C17O (3–2) to estimate the turbulence. C17O has hyperfine lines, but they can be ignored because the hyperfine splitting is minimal ($\sim 0.0012$ km s-1) compared to the spectral resolution ($\sim 0.87$ km s-1). To determine the velocity dispersion, we correct the observed velocity dispersion by deconvolving it from the minor contribution from the velocity resolution (e.g., Huang et al., 2023):
$\sigma_{\rm v}^{2}=\sigma_{\rm obs}^{2}-\frac{\Delta v_{\rm r}^{2}}{8\rm ln2}$ (5)
where $\sigma_{\rm v}$ is the velocity dispersion along the line of sight, $\sigma_{\rm obs}=\Delta V_{\rm FWHM}/(2\sqrt{2\rm ln2})$, and $\Delta v_{\rm r}$ is the spectral resolution of the observations $\sim 0.87\ \rm km\ s^{-1}$. The non-thermal component of the velocity dispersion can be expressed as:
$\sigma_{\rm nth}^{2}=\sigma_{\rm v}^{2}-\sigma_{\rm th}^{2}=\sigma_{\rm v}^{2}-\frac{k_{\rm B}T_{\rm kin}}{m_{\rm obs}}$ (6)
where $\sigma_{\rm nth}$ and $\sigma_{\rm th}$ are the non-thermal and the thermal velocity dispersion, respectively. $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm obs}$ is the mass of the observed molecule (for C17O, $m_{\rm obs}=29\ m_{\rm H}$). $T_{\rm kin}$ is the kinetic temperature of the gas, which we assume to be equal to $\overline{T}_{\rm env}$. Due to the limited velocity resolution, only a small subset of sources had their velocity dispersions resolved in our observation. Kim et al. (2020) reported velocity dispersion values of 0.68, 0.72, 0.65, and 0.70 km s-1, which were traced by the N2H+ line on the same position as HOPS-11, HOPS-164, HOPS-399, and HOPS-403, respectively. These values are comparable to our derived dispersions (see Table LABEL:Tab:parameters), with a difference of $\sim$10%, although their studies focused on larger scales. We utilize the derived velocity dispersions for the following analysis, as our objective is to identify correlations between pairs of parameters, including the non-thermal velocity dispersion. The non-thermal velocity dispersions are listed in column 9 of Table LABEL:Tab:parameters. As shown in Figure 2(G), the median and mean value of velocity dispersion are 0.77 and 0.82 km s-1, respectively.
3.3 Angle Dispersion of B-field
The conventional approach for estimating the angle dispersion of B-fields ($\delta\phi$) is through the measurement of its standard deviation. However, this approach is limited by the 180 ambiguity in polarization angle, which complicates the accurate estimation of angle dispersion. To address this challenge, further methods have been proposed to quantify $\delta\phi$. Heitsch et al. (2001) proposed a variation in which the restriction for large angles was removed. Later Hildebrand et al. (2009) and Houde et al. (2009) applied a more sophisticated technique which uses an analytical derivation of the dispersion in polarization angles within a turbulent molecular cloud. In a recent study, Cortés et al. (2024) proposed an alternative approach to derive the angle dispersion by assuming that the local turbulence perturbing the field follows a Gaussian distribution, and they suggested that 16 Nyquist-sampling segments of B-field are required for the Gaussian distribution under the assumption of 5° for the margin of error and 10° for the mean standard deviation of polarization angles. In this study, we follow the typical method to estimate the angle dispersion $\delta\phi$ by calculating the standard deviation of the distribution of B-field segments. In order to reduce the ambiguity associated with the 180° orientation of the B-field, the distribution of the position angles is shifted when the angles are centred near 0° or 180°. The standard deviation of the shifted distribution is regarded as $\delta\phi$. To avoid the contamination of self-scattering (Huang et al., 2024), the polarization segments within the innermost 400 au region ($\sim 1\arcsec$, corresponding to the beam size) have been masked. We also mask the polarization data for regions beyond a radius of 1200 au, consistent with the scales used to derive the physical parameters in Section 3.1. Protostellar envelopes with less than 16 Nyquist-sampling segments of B-field within 1200 au regions have large uncertainties ($>5^{\circ}$) according to Cortés et al. (2024), and were therefore excluded from our analysis of angle dispersion. Among the remaining 18 protostellar envelopes with sufficient B-field segments, 5 exhibit B-field morphology that is standard-hourglass, 7 are rotated-hourglass, 4 are spiral, and 2 are complex. The angle dispersions $\delta\phi$ for these envelopes are listed in column 10 of Table LABEL:Tab:parameters, and Figure 2(H) shows the distribution of $\delta\phi$ for each type of B-field morphology. Notably, standard-hourglass envelopes tend to have lower angle dispersions. The median and mean values of $\delta\phi$ are 31.9 and 30.1, respectively.
3.4 Fragmentation Level and Disk Properties
[…]
Figure 4: The stacked histogram of disk parameters for protostars with different B-field structure. Upper row: disk mass $M_{\rm disk}$ (panel A), disk radius $R_{\rm disk}$ (B). Bottom row: disk inclination angle $\theta_{\rm Inc}$ (panel C and D). Dashed and solid lines indicate the mean and median values of the distribution. In panel D, “Face-on”, “Inclined”, and “Edge-on” indicate the inclination angle within a range of 0–30, 30–60, and 60–90, respectively. The disk mass, disk radius, and disk inclination angle are obtained from Tobin et al. (2020).
Tobin et al. (2020) identified protostellar disks in the Orion molecular clouds using both VLA and ALMA observations, with further refinements presented in their subsequent study of Tobin et al. (2022). All objects in our sample are included within the VANDAM survey, which enables us to estimate the multiplicity of these systems. Palau et al. (2013) defined the fragmentation level as the number of millimeter sources inside a region with a radius of 0.05 pc for massive dense cores, whereas Busquet et al. (2016) used a different radius of 0.15 pc. We follow the above methods, but define a different radius of 1200 au, which is the scale for estimating the physical parameters above. Then the number of fragments (including small-scale fragmentation, that was only resolved in the survey with higher resolution of $\sim$40–400 au-scale companions) identified from the VANDAM survey within this innermost envelope region is indicative of the fragmentation level $N_{\rm F}$. It is possible that the outer ($>1200$ au) regions may influence the fragmentation of the envelopes, given that the typical scale for the star-forming envelopes is up to several thousand au (e.g., Myers et al., 2000). We thus define larger radii of 1800 au and 2400 au for estimating extra contribution of $N_{\rm F}$, but with less weighting values of 0.5 and 0.25, respectively. As shown in Figure 3, which shows the 0.87 mm dust emission map of an example HOPS-78, a protostar (black asterisk) is located in the innermost region with a radius of 1200 au (black dashed circle), then $N_{\rm F}$ in this region is counted as 1. While there are 2 (blue asterisks) and 1 (red asterisk) protostars in annular regions with inner and outer radii of 1200 and 1800 au (blue dashed circle), and 1800 and 2400 au (red dashed circle), then the $N_{\rm F}$ for each region is counted as 1 and 0.25, respectively. The total $N_{\rm F}$ for HOPS-78 is therefore assigned a value of 2.25. Column 11 of Table LABEL:Tab:parameters lists the fragmentation level for the sample.
The mass, radius, and inclination angle of the disks listed in column 12–14 of Table LABEL:Tab:parameters are also taken from Tobin et al. (2020). The upper row of Figure 4 shows the distribution of disk mass and radius, with median and mean values of 118.5 and 171.7 M, and 69.7 and 79.2 au, respectively. While the bottom row of Figure 4 presents the distribution of disk inclination angle, which is approximated as $\theta_{\rm inc}=\arccos(\Theta_{\rm min}/\Theta_{\rm maj})$, where $\Theta_{\rm maj}$ and $\Theta_{\rm min}$ are the deconvolved major and minor axes of the disk. Here the face-on and edge-on configurations are defined with an inclination angle of $\theta_{\rm Inc}$ = 0, and $\theta_{\rm Inc}$ = 90, respectively. The mean and median values of the inclination angle are 44.3 and 45.9, respectively. Note that the finite resolution and vertical thickness of the disks (Sheehan et al., 2022; Lin et al., 2023; van’t Hoff et al., 2023) will bias this simple inclination estimate against 90, which is probably why it peaks at $\sim 45^{\circ}$ and not 60.
4 Discussion
In Section 3, we have derived the physical parameters of dust temperature, envelope gas mass, gas column density, and gas number density for the entire sample of BOPS within the inner region of 6″ (corresponding to a scale of 2400 au, with a radius of 1200 au). The distributions of these parameters are shown in Figure 2. We find from Figure 2(A–B) that protostars with high bolometric luminosity ($>$ 50 L) and/or envelope dust temperature ($>$ 25 K) tend to exhibit spiral or rotated-hourglass B-field morphology. In Figure 2(C–F), sources with compact polarized emission (self-scattering) tend to have relatively lower total flux density, envelope masses, gas column densities, and gas number densities, whereas protostars exhibiting a standard-hourglass B-field pattern tend to exhibit relatively larger values in these parameters. However, these standard-hourglass protostars have relatively low velocity dispersion, which is potentially indicative of the magnetic braking effect, whereby the strong field exerts a braking influence on the rotational rate (as shown in panel G of Figure 2). In Figure 2(H), we find standard-hourglass-B-field protostars have relatively low angle dispersion, in contrast to rotated-hourglass-, spiral-, and complex-field sources, which display a relatively larger angle dispersion.
The uncertainties of the derived physical parameters are due to unknowns and errors in factors such as dust opacities, dust temperatures, and observed flux. The uncertainty in physical parameters resulting from dust opacity, an internally consistent systematic error, does not significantly impede the identification of correlations between physical parameters, as these are calculated in a uniform manner, which translates the distributions up or down. Whereas the dust temperature and the flux calibration are considered random. For the temperature, based on the model that we used, we consider typical uncertainties of 20–30%, and we expect flux calibration to be accurate within 10%. All combined, we derive uncertainties of about 30–40% for the masses and densities.
The disk properties are shown in Figure 4. We find from the first row that, the distribution of disk mass and radius appears random across different types of B-field morphologies. In the bottom row, in both types of sources dominated by self-scattering and exhibiting rotated-hourglass morphologies, with the mean polarization angle preferentially aligned parallel to the outflow direction, the disk inclination angle is found to be either very inclined or edge-on. However, in standard-hourglass sources, with the mean polarization angle oriented perpendicularly to the outflow, the disk inclination angle is typically found to be face-on. This is probably due to the fact that some disks in the VANDAM project are potentially influenced by the envelope emission (HOPS-87, HOPS-317S, and HOPS-403, see Tobin et al., 2020).
To statistically examine the relations between two pairs of parameters, we performed both the Pearson linear correlation coefficient ($r_{p}$) and Spearman’s rank correlation coefficient ($r_{s}$). The value of $r_{p}$ and that of $r_{s}$ are each bounded from -1 to 1, for negative to positive slope correlations, respectively. In each case, the closer the absolute value is to 1, the stronger the correlation. The p–value is used for the purpose of testing the null hypothesis that the two variables are uncorrelated. In general, statistically significant correlations are identified by p–value $<$ 0.05. Here we analyze the following figures, and based on these details, we will motivate a more detailed analysis on fragmentation and the possible effect of magnetic braking on disk formation.
4.1 Physical Envelope Properties
Figure 5: Left panels: Distributions between the bolometric luminosity $L_{\rm bol}$ and the envelope gas mass $M_{\rm env}$ (panel A), the gas column density $N_{\rm gas}$ (C), the gas volumn density $n_{\rm gas}$ (E), and the non-thermal velocity dispersion $\sigma_{\rm nth}$ (G). Right panels: Distribution between the mean dust temperature $\overline{T}_{\rm env}$ and the envelope gas mass $M_{\rm env}$ (B), the gas column density $N_{\rm gas}$ (D), the gas volume density $n_{\rm gas}$ (F), and the non-thermal velocity dispersion $\sigma_{\rm nth}$ (H).
The dust temperature is expected to be correlated with the bolometric luminosity if the dust is primarily heated by the stellar radiation (see Section 3.1). During the early stages of star formation (Class 0 and early Class I), the protostar is deeply embedded within a dense envelope. In this phase, the luminosity is predominantly due to the accretion of material from the envelope through the disk onto the protostar, which releases gravitational energy. High-density regions are often associated with intense accretion activity as the protostar accumulates material. The inflow of material from the envelope onto the protostar leads to an increase in $L_{\rm bol}$. Additionally, a more massive envelope absorbs and re-radiates more energy and it is more efficient at cooling due to its higher optical depth. We plot the distribution between these physical parameters, as shown in Figure 5(A–F). However, our results reveal no significant correlations between $L_{\rm bol}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, and between $\overline{T}_{\rm env}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, as their correlation coefficients are close to 0, with p–values much higher than 0.05. The lack of a direct correlations are not surprising due to the dynamic nature of protostellar evolution. Protostellar envelope masses are likely influenced by both the final stellar mass of the protostar and the evolutionary stage, with envelope masses decreasing rapidly over time (Fischer et al., 2017). Luminosity, on the other hand, is influenced by episodic accretion, which may vary independently of the envelope mass (e.g., Zakri et al., 2022; Narang et al., 2024).
Non-thermal motions, often driven by turbulence, are of significant importance in star-forming regions and affect the overall dynamics and evolution of the protostellar system. Therefore, $\sigma_{\rm nth}$ serves as an indicator of turbulence level, which enhance the accretion process by increasing the rate at which material is fed onto the protostar. As shown in Figure 5(G–H), a slight positive correlation is observed between $\sigma_{\rm nth}$ and $L_{\rm bol}$, as well as between $\overline{T}_{\rm env}$ and $L_{\rm bol}$. Protostars with higher value of $L_{\rm bol}$ typically drive powerful outflows and winds, injecting energy into the surrounding medium and thereby increasing the non-thermal velocity dispersion. In high turbulent regions, the non-thermal velocity dispersion provide support against gravitational collapse. Once, however, collapse is triggered, the turbulent motions channel material onto the forming protostar, influencing its luminosity. Theoretical models also suggest that higher levels of turbulence ($\sigma_{\rm nth}$) lead to more efficient accretion processes, leading to increased protostellar luminosities (e.g., Krumholz et al., 2005; Padoan et al., 2011; Federrath, 2015). Additionally, in regions with high $\sigma_{\rm nth}$, which are characterised by intense star formation activity (e.g., jets and winds), one might expect that higher dust temperatures may result from the combined effects of shock heating and enhanced radiative feedback from very young protostars. This is due to the fact that the surrounding dust is heated by protostellar outflows, winds and radiation.
We note that these results will be influenced by the assumption of a typical temperature of 20 K. However, we find that 9 protostars exhibit relatively high values of $L_{\rm bol}$ and $\overline{T}_{\rm env}$. Of these, 3 display a rotated hourglass field shape (HOPS-168, HOPS-288, and HOPS-370), 5 of the protostars are spiral (HOPS-124, HOPS-182, HOPS-361N, HOPS-361S, and HOPS-384), and one protostar with compact polarized emission (HOPS-84).
4.2 Fragmentation
Figure 6(A–C) show the relation between $N_{\rm F}$ and $M_{\rm env}$, $\sigma_{\rm nth}$, and $\delta\phi$, respectively. It is important to note that the values of $N_{\rm F}$ are dominated by multiples within the inner 2400 au region, a scale that is consistent with the one used to estimate the physical parameters as discussed in Section 3. For protostars located in multiple systems with separation distances of less than 1200 au, the fragmentation level correlations are plotted only once, using the protostar with the highest dust emission as the reference point. Combined with the statistical test methods, we discuss the potential influence of specific factors on envelope fragmentation at the 2400 au scale in the following.
Optical and near infrared studies suggest that the multiplicity rate drops with decreasing mass (e.g., Chini et al., 2012, 2013). We investigated the correlation between $N_{\rm F}$ and $M_{\rm env}$, and the result is shown in Figure 6(A). Based on the correlation coefficients shown in Figure 6(A), there is no clear correlation between $N_{\rm F}$ and $M_{\rm env}$. This result may suggest that the correlation observed in the optical/infrared is not yet established at the very early stages of star and cluster formation. It is possible, therefore, that the final number of members in a multiple system is determined after multiple star formation episodes, as suggested by both theoretical (e.g., Myers, 2011) and observational(e.g., Bik et al., 2012) studies. Furthermore, the process of fragmentation and the formation of multiple objects are typically governed by a complex interplay of factors beyond just the envelope mass. These interactions may lead to significant variability in outcomes, masking any simple correlation between envelope mass and multiplicity. Previous observational studies of massive dense cores suggest there is no correlation between mass and multiplicity within a radius of 0.05 pc of the dust peak (Palau et al., 2013).
In recent decades, several studies on the turbulence scenario (e.g., Padoan et al., 2001; Vázquez-Semadeni et al., 2007) have demonstrated that turbulence is capable of generating density structures through the action of supersonic shocks, which compress and fragment the gas efficiently, particularly in the presence of significant turbulence. Subsequent studies have indicated that different types of turbulence, i.e. compressive and solenoidal, exert divergent effects on fragmentation (e.g., Federrath et al., 2008). Compressive turbulence is conducive to fragmentation, whereas solenoidal turbulence has the effect on slowing down fragmentation (e.g., Federrath et al., 2010; Girichidis et al., 2011). As shown in Figure 6(B), the fragmentation level $N_{\rm F}$ is weakly correlated with $\sigma_{\rm nth}$, and this correlation becomes more evident when the protostar HOPS-361S with $N_{\rm F}=4$ is excluded, with correlation coefficients of $r_{p}\sim 0.42$, $r_{s}\sim 0.40$, p–value $\ll 0.01$. This positive correlation between $N_{\rm F}$ and $\sigma_{\rm nth}$ in our result suggests that turbulence tends to promote the fragmentation, indicating that compressive turbulence may be more dominant on envelope scales. However, one should note that $\sigma_{\rm nth}$ may also include contributions from rotational and infalling motions on envelope scales, which cannot be distinguished with the limited spectral resolution of C17O. Therefore, deeper studies with sufficient spectral resolution should be carried out to reach more robust conclusions.
On the other hand, the fragmentation level in massive dense cores could be explained by a B-field/turbulence balance (Palau et al., 2013) or changes in the mass-to-flux ratio (Busquet et al., 2016; Añez-López et al., 2020; Palau et al., 2021) on scales of several thousands au, suggesting a potential correlation between B-fields and fragmentation on envelope scales. One of the most longstanding methods for estimating the B-field strength is the Davis-Chandrasekhar-Fermi (DCF, Davis, 1999; Chandrasekhar et al., 1953) method, which relies on the principle that turbulent motions should visibly affect the distribution of the B-field in the plane of the sky. The basic form of the DCF formula is $B_{\rm{pos}}=\sqrt{4\pi\rho}\sigma_{\rm nth}/\delta\phi$, where $\rho$ is the gas density, and $\delta\phi$ is the measured angular dispersion in B-field position angles (as discussed in Section 3.3). Since $\sigma_{\rm nth}$ is barely resolved for most sources in our sample, which significantly affect the estimation of B-field strength. However, the above equation indicates that the field strength is inversely proportional to the dispersion in the polarization position angles, which provides a potential way to indirectly study the correlation between field strength and other physical parameters on envelope scales. In order to statistically test this correlation, we plot the distribution between the angle dispersion of the B-field $\delta\phi$ and the fragmentation level $N_{\rm F}$, as shown in Figure 6(C). Combing correlation coefficients, we find that $N_{\rm F}$ is positive correlated with $\delta\phi$. Even when excluding the outlier point of protostar HOPS-361S with $N_{\rm F}=4$, the positive correlation ($r_{p}\sim 0.50$, $r_{s}\sim 0.55$, p–value $\sim 0.03$) remains significant. This suggests that higher angle dispersion in the B-field is associated with higher multiplicity, and thus, the fragmentation tends to be suppressed by B-field. Previous studies have indicated that strong B-fields suppress fragmentation, as they constitute an additional form of support against gravitational contraction and also via the magnetic effect (inward pinching and gathering of the field lines dragged during collapse makes the magnetic strength larger at inner radius, and acts against rotation), which reduces the amount of angular momentum in the inner part of cores (e.g., Galli et al., 2006; Mellon et al., 2008; Hennebelle et al., 2008; Price et al., 2009; Myers et al., 2014; Cunningham et al., 2018; Kwon et al., 2019; Yen et al., 2021).
In addition, multiplicity may also decrease with evolution from Class 0 to Class I (Tobin et al., 2020, 2022). However, since the BOPS sample predominantly consists of objects at the Class 0 stage, the sample size in our study is insufficient to differentiate evolutionary factors over time.
Figure 6: Panel A–C: Distributions between the fragmentation level $N_{\rm F}$ and the envelope mass $M_{\rm env}$ (panel A), the non-thermal velocity dispersion $\sigma_{\rm nth}$ (panel B), and the angle dispersion of B-field $\delta\phi$ (panel C). Panel D–F: Distributions between the disk radius $R_{\rm disk}$ and the alignment of the B-field and outflow axes $|\Delta\theta|$ (panel D), and between the disk mass $M_{\rm disk}$ and $|\Delta\theta|$ (panel E and F). In panel F, disk masses with their disk radii less than 100 au are derived from Sheehan et al. (2022).
4.3 B-field and Disk Formation
Observations indicate that cloud cores in star-forming environments are magnetized (e.g., Crutcher, 1999). Theoretical models suggest that B-field tension forces deflect infalling gas towards the equatorial plane, resulting in a dense, flattened pseudo-disk (e.g., Allen et al., 2003). During the collapse phase, the field lines are drawn inward into a pinched geometry (e.g., Girart et al., 2006, 2009; Stephens et al., 2013; Kwon et al., 2019), which strengthens the B-field and carries the angular momentum outward. This process is known as magnetic braking, which reduce angular momentum by one order of magnitude, potentially suppressing the formation of a centrifugally supported disk (e.g., Galli et al., 2006). However, some studies propose that an initial misalignment between the B-field and the rotation axis would lead to weaker magnetic braking of the collapsing core, enabling earlier disk formation (e.g., Joos et al., 2012; Li et al., 2013; Hirano et al., 2019; Machida et al., 2020). Recent observational studies have also reported that a disk can be formed even in young stellar objects with B-fields aligned with the rotation axis (Kwon et al., 2019; Yen et al., 2021).
Huang et al. (2024) have revealed the dust polarized continuum emission and outflow signatures in protostars, finding that 40% of the protostars exhibit a mean B-field direction that is approximately perpendicular to the outflow on envelope scales. However, the relative orientation in the remaining sample appears to be random. The disk mass and disk radius taken from Tobin et al. (2020) are shown in the first row of Figure 4, which are both randomly distributed. We combine the above disk properties and envelope misalignment between the B-field and outflow, as shown in Figure 6(D–F). Protostars with complex and unresolved polarized emission, as well as those with disks significantly affected by envelope emission (HOPS-87, HOPS-317S, and HOPS-403), are excluded in our analysis. This exclusion is due to the substantial uncertainty in the averaged B-field direction or disk radius/mass for these sources. In Figure 6(D), protostars with a standard hourglass or rotated hourglass B-field structure exhibit a positive correlation between disk radius $R_{\rm disk}$ and the misalignment angle between magnetic and outflow axes $|\Delta\theta|$. In contrast, protostars with spiral field structure exhibits a random distribution. When accounting for envelopes with various B-field morphologies, a positive correlation is significant in the distribution (as shown in panel D of Figure 6). Similar discernible trends for each type of B-field morphology can be found in Figure 6(E), which illustrate a distribution between the disk mass $M_{\rm d}$ and $|\Delta\theta|$. A weaker positive correlation is noted between $M_{\rm disk}$ and $|\Delta\theta|$ when including all envelopes (standard hourglass, rotated hourglass, and spiral). This reduced correlation compared to $R_{\rm disk}$$|\Delta\theta|$ relation suggest that the derived disk mass from Tobin et al. (2020) may have potential uncertainties.
Sheehan et al. (2022) found that the small disks (with radius less than $\sim 100$ au) often have their masses overestimated compared to radiative transfer model, with the discrepancy becoming more pronounced for lower-mass disks. To address this, they proposed a log-linear relation to reconcile the disk masses derived from Tobin et al. (2020) with those obtained through radiative transfer modelling, expressed as:
$\log(M_{\rm disk}^{\rm Sheehan})=\frac{\log(M_{\rm disk})-(0.83\pm 0.05)}{(0.67\pm 0.03)}$ (7)
Figure 6(F) shows the distribution between the re-estimated disk mass and $|\Delta\theta|$. The positive correlation between these variables remains apparent, suggesting that the uncertainty in disk mass derived from Tobin et al. (2020) does not significantly affect the observed correlation.
If the bottom-left point in Figure 6(D–F) is excluded, the updated correlation coefficients are $r_{p}\sim 0.54$, $r_{s}\sim 0.59$, p–value $<0.01$ for panel A, $r_{p}\sim 0.49$, $r_{s}\sim 0.53$, p–value $\sim 0.02$ for panel B, and $r_{p}\sim 0.30$, $r_{s}\sim 0.41$, p–value $\sim 0.05$, respectively. These results indicate that the correlations between disk properties and the misalignment angle remain consistent, even when the bottom-left point is excluded, suggesting that the correlation is less importantly affected by this point. The p-values, being $\lesssim$ 0.05, suggest that these correlations are statistically significant.
Huang et al. (2024) applied both the intensity-weighted method (which gives more weight to the inner protostellar envelope) and the error-weighted method (which provides an averaged weight across the selected envelope) to estimate the mean direction of the (B-field). The mean direction distributions of B-field obtained from both weighting methods were found to be comparable across the entire sample, suggesting that the measured B-field is minimally influenced by gravitational collapse and outflows. In this case, the measured B-field is likely representative of the original pre-stellar core B-field. Furthermore, the outflow direction aligns with the angular momentum of the original core. Our findings suggest that a greater misalignment between these two directions lead to weaker magnetic braking of the collapsing core and result in a more prominent disk, which is consistent with the misalignment-induced disk formation model (e.g., Joos et al., 2012).
5 Conclusions
We have calculated and analyzed the physical parameters for 55 protostelar envelopes within the BOPS sample. Non-thermal velocity dispersion is often associated with turbulence, and the polarized dust emission is thought to trace B-fields by magnetically aligned grains on envelope scales. The combination of envelope properties from the BOPS survey (Huang et al., 2024) and the disk properties from the VANDAM survey (Tobin et al., 2020) allows us to link the turbulence and B-field in the envelope to the resulting fragmentation and disk properties. The main results are as follows:
  1. 1.
    Protostellar envelopes with compact polarized dust emission tend to have a lower flux density, envelope mass, gas column density, and gas number density. While envelopes with a standard-hourglass field shape show higher values for these parameters.
  2. 2.
    Envelopes with a standard hourglass B-field structure generally show relatively low non-thermal velocity dispersion. In contrast, envelopes with a spiral B-field structure tend to have higher non-thermal velocity dispersion, indicating a more turbulent environment.
  3. 3.
    The disk mass and disk radius show no significant correlation with the different B-field structures in the envelopes. The disk inclination angles might affect the measurement of the observed B-field morphology, but the finite resolution and vertical thickness of the disks could introduce biases in the measurement of inclination angles.
  4. 4.
    The dust temperature and bolometric luminosity are not correlated with envelope masses and densities, despite these quantities being derived from luminosity. However, protostars with a significant velocity dispersion or high turbulence levels tend to exhibit increased luminosity and higher dust temperature. This correlation is likely attributed to enhanced accretion rates and energetic feedback processes, suggesting that the luminosity is predominantly driven by accretion.
  5. 5.
    The lack of correlation between fragmentation level and envelope mass is possibly due to the complex interplay of various factors influencing fragmentation beyond just the envelope mass. The fragmentation level is weakly correlated with the non-thermal velocity dispersion, suggesting that turbulence may promote envelope fragmentation on 2400 au scales. However, the non-thermal velocity dispersion may include contributions from disk rotation, envelope rotation, and/or infalling motions, necessitating further observations to confirm this scenario. Additionally, the angle dispersion of the B-field, which is inversely proportional to the field strength, is positively correlated with the fragmentation level, indicating that the pinched field suppresses the fragmentation, as it provides additional support against gravitational collapse, and/or brakes the rotation efficiently.
  6. 6.
    A more significant misalignment between the magnetic and outflow axes is associated with larger disk radii and disk masses. This suggests that increased misalignment results in weaker magnetic braking during core collapse, which aligns with the misalignment-induced disk formation model.
It is important to note that our results are based on the limited spectral resolution of C17O observations, which may influence the conclusions drawn with respect to the velocity dispersion. Further observations with higher sensitivity and resolution are needed to confirm these findings and refine out conclusions.
B.H. and J.M.G. acknowledge support by the grant PID2020-117710GB-I00 (MCI-AEI-FEDER, UE). B.H. also acknowledges financial support from the China Scholarship Council (CSC) under grant No. 202006660008. This work is also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. M.F.L. acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). M.F.L. also acknowledges the warmth and hospitality of the ICE-UB group of star formation. L.W.L. acknowledges support by NSF AST-1910364 and NSF AST-2307844. Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. W.K. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00342488). The authors acknowledge Anaëlle Maury for helpful discussions, Jacob Labonte for early analysis of the BOPS data. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00086. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.