### v2

# The protostars in Orion: Characterizing the properties of their magnetized envelopes

Abstract
We present a study connecting the physical properties of protostellar envelopes to the morphology of the envelope-scale magnetic field. We used the ALMA polarization observations of 61 young prtostars at 0.87 mm on $\sim 400-3000$ au scales from the B-field Orion Protostellar Survey to infer the envelope-scale magnetic field, and used the dust emission to measure the envelope properties on comparable scales. We find that protostars showing standard-hourglass-field morphology tend to have larger masses and lower velocity dispersions in their envelopes, whereas systems with spiral-field morphologies have higher velocity dispersion. Combining with the disk properties taken from the Orion VLA/ALMA Nascent Disk and Multiplicity survey, we connect envelope properties to fragmentation. Our results show that the fragmentation level is positively correlated with the angle dispersion of the magnetic field, suggesting that the envelope fragmentation tends to be suppressed by the magnetic field. We also find that protostars exhibiting standard hourglass magnetic field structure have a smaller disk and smaller angle dispersion of the magnetic field than other field configurations, specially the rotated hourglass, but also the spiral and others, suggesting a more effective magnetic braking in the standard hourglass morphology of magnetic fields. Nevertheless, significant misalignment between the magnetic field and outflow axes tends to reduce magnetic braking, leading to the formation of larger disks.
Star formation - star forming regions - magnetic fields - interstellar magnetic fields - circumstellar envelopes - circumstellar disks
1 Introduction
Fragmentation is believed to be a critical process in the formation of star and stellar systems (e.g., Matsumoto et al., 2003). Initially, turbulence induces chaotic motions and density fluctuations across various scales, compressing the gas in certain regions and leading to localized density enhancements (e.g., Mac Low et al., 2004). Within these dense regions, gravitational forces overcome thermal pressure if the mass exceeds the Jeans mass. As the cloud collapses, the large structures fragment into smaller, dense cores, each of which forms individual stars or multiple star systems (e.g., McKee et al., 2007). This hierarchical fragmentation continues until individual cores reach densities high enough to resist further fragmentation (e.g., Beuther et al., 2015). However, the specific mechanisms by which a cloud core fragments to ultimately form a multiple system or small cluster remain uncertain.
Theoretical and numerical studies (e.g., Maury et al., 2022; Offner et al., 2023) suggest that cores with some angular momentum are prone to fragmentation into multiple objects, indicating the need for additional mechanisms to reduce the fragmentation to the observed levels. Since non-ideal MHD effects do not inhibit protostellar multiplicity (e.g., Wurster et al., 2019), further research points to turbulence and magnetic fields (hereafter, B-fields) as playing a pivotal role in the fragmentation process (e.g., Padoan et al., 2001; Price et al., 2009). For instance, compressive turbulence promotes fragmentation and the formation of structures with high density contrasts, whereas solenoidal turbulence (e.g., Federrath et al., 2010; Girichidis et al., 2011) suppresses fragmentation and maintains a smoother density structure within the cloud. On the other hand, the ratio of rotational to magnetic energy has been found to influence core fragmentation (e.g., Machida et al., 2008). Moreover, some studies suggest that B-field tends to suppress fragmentation by providing additional support against gravitational collapse, and/or through magnetic braking, which carries away the angular momentum from the inner regions of the core (e.g., Commerçon et al., 2011; Guszejnov et al., 2021; Mignon-Risse et al., 2021).
Dust polarization is the most commonly used tracer of B-fields in star-forming regions, as “radiative torques” (e.g., Hoang et al., 2009; Andersson et al., 2015) tend to align spinning and elongated dust grains with their long axes perpendicular to the ambient B-field direction. Over the last decades, dust polarization observations carried out with (sub)millimeter interferometers have increasingly proven effective in mapping B-fields at the scales of cores ($\sim 10^{4}$ au) and envelopes ($\sim 10^{2}$ to $10^{4}$ au) (e.g. Girart et al., 1999, 2013; Cox et al., 2018; Galametz et al., 2018; Hull et al., 2019; Le Gouellec et al., 2020; Cortés et al., 2021). The theory predicts that the B-field regulates the collapse of clouds, whereas ambipolar diffusion allows the formation of “supercritical” dense cores, where the gravitational force dominates over the magnetic support. As cloud cores collapse, B-field lines are dragged inward to be drawn into pinched morphology. This has been observed in various studies (e.g., Girart et al., 2006, 2009; Qiu et al., 2014; Beltrán et al., 2019; Le Gouellec et al., 2019; Hull et al., 2020). In the case of flux freezing, the pinched geometry in the pseudo-disk strengthens the B-field and carries away angular momentum, leading to catastrophic magnetic braking and potentially preventing the disk formation (e.g., Galli et al., 2006). However, disk may form if the magnetic and rotational axes are misaligned (e.g., Joos et al., 2012; Li et al., 2013; Machida et al., 2020). To statistically test these scenarios related to fragmentation and star/disk formation, it is crucial to understand the connection between envelope properties and the resulting fragmentation.
Recently, the first results from the B-fields Orion protostellar survey (BOPS, with a resolution of $\sim 1^{\prime\prime}$, corresponding to $\sim$ 400 au, Huang et al., 2024) presented the largest polarization study of B-fields in protostellar envelopes using dust polarization observations of 61 protostars in Orion. We have successfully mapped the B-fields morphology at scales ranging from several hundred to a few thousand au, and classified the protostars using two criteria: polarization/B-field structure and the alignment between the outflow directions and velocity gradients. On one hand, the morphological categories of the B-field for protostars are classified as: “standard hourglass” (with the main axis coincident with the outflow axis), “rotated hourglass” (similar to the “standard hourglass” but rotated by $\sim 90^{\circ}$, with its axis perpendicular to the outflow), “spiral” configuration, “complex” field morphology, “not enough data” (insufficient polarized data to identify the B-field structure), and “self-scattering” (compact polarized emission in the innermost regions, with the polarization direction perpendicular to the disk major axis). On the other hand, BOPS protostars are classified as “Perp-Type” (velocity gradient is perpendicular to the outflow), “Nonperp-Type” (velocity gradient is not perpendicular to the outflow), and “Unres-Type” (velocity gradient is unresolved). In addition, observations from a major Orion interferometric study, known as the Orion VLA/ALMA Nascent Disk and Multiplicity (VANDAM, Tobin et al., 2020), have identified protostellar disks in the Orion molecular clouds with a resolution of $\sim 0\farcs 1$ (corresponding to $\sim$ 40 au). This study provided data on the radius and mass of these disks. Since all BOPS protostars were originally observed at high resolution in the VANDAM survey, the combination of these datasets enables a comprehensive test of the magnetic braking model and allows for a more effective linkage of magnetic and turbulent properties to the characteristics of protostellar disks and multiplicity.
This paper builds upon the first results of BOPS (Huang et al., 2024) and is organized as follows: Section 2 introduces the observations and data reduction processes. Section 3 presents the methods for the calculation and estimation of physical parameters, the velocity dispersion, the angle dispersion of the B-field, and the fragmentation level. A detailed discussion of the correlation between these parameters is provided in Section 4. Finally, the conclusions are summarized in Section 5.
2 Observations
The 870 $\mu$m (Band 7) dust polarization observations were taken with ALMA, and the spatial resolution and the Largest Angular Scale (LAS) are $\sim$1″and $\sim$8″, corresponding to a linear resolution of $\sim$400 and $\sim$3200 au at a distance of $\sim$400 pc (see Huang et al. 2024 for further details on the observation setup).
[…]
For the C17O (3–2) line data, we used the CASA uvcontsub task to perform continuum fitting and subtraction in the $u$-$v$ plane. We then utilized the tclean task to produce Stokes I image cubes using self-calibrated, continuum-subtracted data for each spectral window. The C17O (3–2) line, optically thin in dense regions (with spectral resolution of $\sim$ 0.87 km s-1, Huang et al., 2024), was employed to trace the velocity field of the envelope.
3 Analysis
The first results of the BOPS have revealed the dust polarized continuum emission for 61 protostars (Huang et al., 2024), and some parameters are listed in Appendix A. In this work, we use the classification criteria based on the B-field/polarization structure because the velocity gradients in a large fraction of the sample are unresolved. Then the 61 protostars are classified into 5 types: “standard hourglass”, “rotated hourglass”, “spiral”, “others” (including “complex” and “not enough data”), and “self-scattering”, as shown in Appendix B. The following sections presents the estimation of physical envelope parameters, velocity dispersion, the angle dispersion of B-field, and fragmentation level for all BOPS protostars. These properties will then be compared to the disk properties to understand which factors could potentially influence the envelope fragmentation and whether the initial misalignment between the B-field and outflows could result in reduced magnetic braking during disk formation.
3.1 Physical Envelope Properties
[…]
$M_{\rm dust}=\frac{S_{\nu}d^{2}}{\kappa_{\nu}B_{\nu}(T_{\rm dust})}$ (1)
where $d$ is the distance, $\nu$ is the frequency corresponding to the observed wavelength, $B_{\nu}(T_{\rm dust})$ is the Planck function at the dust temperature of $T_{\rm dust}$, and $\kappa_{\nu}$ is the dust opacity. We adopt $\kappa_{870\mu{\rm m}}\approx 1.84\ {\rm cm}^{2}\ {\rm g}^{-1}$, which is appropriate for dusty particles with ice mantles at densities of approximately $10^{6}$ cm-3, as reported by Ossenkopf et al. (1994).
Table 1: Envelope and disk parameters for the BOPS sample\$$\dagger$\$$\dagger$\$footnotemark:
Name $L_{\rm bol}$ $S_{\nu}$ $\overline{T}_{\rm env}$ $T_{\rm c}$ $M_{\rm env}$ $N_{\rm gas}$ $n_{\rm gas}$ $\sigma_{\rm nth}$ $\delta\phi$ $N_{F}$ $N_{B}$
[…]
Standard
HOPS-11 9.00 0.40 26.1 17.7 $\pm$ 0.9 0.26 1.3E+23 5.4E+06 0.67 $\pm$ 0.20 / 1.00 6
HOPS-87 36.49 3.89 37.0 21.2 $\pm$ 1.9 1.46 7.3E+23 3.0E+07 0.43 $\pm$ 0.13 17.6 $\pm$ 1.4 1.00 86
HOPS-359 10.00 0.91 26.8 17.4 $\pm$ 2.0 0.62 3.1E+23 1.3E+07 0.45 $\pm$ 0.11 29.2 $\pm$ 2.3 1.00 80
HOPS-395 0.50 0.58 20.2 14.3 $\pm$ 1.1 0.49 2.4E+23 1.0E+07 0.40 $\pm$ 0.12 15.0 $\pm$ 2.0 2.00 29
HOPS-400 2.94 1.18 21.5 15.8 $\pm$ 0.7 1.01 5.0E+23 2.1E+07 0.45 $\pm$ 0.22 18.2 $\pm$ 1.9 2.00 48
HOPS-407 0.71 0.61 20.2 17.8 $\pm$ 2.0 0.59 2.9E+23 1.2E+07 0.44 $\pm$ 0.12 10.5 $\pm$ 1.0 1.50 53
OMC1N-8-N / 0.63 / 37.5 $\pm$ 10.5 0.31 1.6E+23 6.5E+06 0.43 $\pm$ 0.31 20.6 $\pm$ 2.1 2.00 51
Rotated
HH270IRS 7.70 0.49 25.1 16.7 $\pm$ 0.1 0.35 1.7E+23 7.2E+06 0.45 $\pm$ 0.15 33.8 $\pm$ 4.1 2.00 35
HOPS-78 8.93 0.81 26.0 22.6 $\pm$ 2.1 0.48 2.4E+23 9.9E+06 0.59 $\pm$ 0.27 40.2 $\pm$ 4.3 2.25 45
HOPS-168 48.07 0.51 39.6 16.5 $\pm$ 1.4 0.17 8.4E+22 3.5E+06 0.64 $\pm$ 0.26 28.4 $\pm$ 3.9 2.25 28
HOPS-169 3.91 0.67 22.2 19.2 $\pm$ 1.4 0.47 2.3E+23 9.7E+06 0.30 $\pm$ 0.18 32.9 $\pm$ 3.9 1.00 37
HOPS-288 135.47 1.60 51.4 13.7 $\pm$ 1.0 0.43 2.1E+23 8.9E+06 1.34 $\pm$ 0.51 24.5 $\pm$ 2.9 3.00 37
HOPS-310 13.83 0.58 29.1 17.0 $\pm$ 1.2 0.33 1.6E+23 6.8E+06 0.75 $\pm$ 0.15 / 1.00 9
HOPS-317S 4.90 2.33 22.9 14.7 $\pm$ 0.9 1.95 9.7E+23 4.0E+07 0.45 $\pm$ 0.16 / 1.25 15
HOPS-370 360.86 1.08 65.6 25.1 $\pm$ 1.9 0.21 1.0E+23 4.3E+06 0.80 $\pm$ 0.32 33.8 $\pm$ 3.1 1.00 62
HOPS-401 0.61 0.37 20.2 14.8 $\pm$ 1.0 0.37 1.9E+23 7.7E+06 0.69 $\pm$ 0.12 / 1.00 13
HOPS-409 8.18 0.38 25.4 30.8 $\pm$ 7.9 0.23 1.2E+23 4.8E+06 0.83 $\pm$ 0.34 17.0 $\pm$ 2.5 1.25 25
OMC1N-4-5-ES / 0.72 / 36.1 $\pm$ 8.9 0.28 1.4E+23 5.8E+06 0.63 $\pm$ 0.21 34.6 $\pm$ 4.1 2.25 66
Spiral
HOPS-96 6.19 0.44 23.8 25.8 $\pm$ 4.3 0.29 1.4E+23 6.0E+06 0.86 $\pm$ 0.15 / 1.00 11
HOPS-124 58.29 1.43 41.6 20.3 $\pm$ 0.6 0.48 2.4E+23 9.9E+06 1.20 $\pm$ 0.43 / 2.25 8
HOPS-182 71.12 1.43 43.7 16.9 $\pm$ 1.2 0.42 2.1E+23 8.7E+06 1.44 $\pm$ 0.54 46.4 $\pm$ 4.6 3.50 51
HOPS-303 1.49 0.42 20.6 21.4 $\pm$ 1.0 0.37 1.9E+23 7.8E+06 0.56 $\pm$ 0.21 / 1.00 13
HOPS-361N 478.99 1.70 70.4 18.6 $\pm$ 3.1 0.36 1.8E+23 7.5E+06 0.69 $\pm$ 0.33 44.2 $\pm$ 3.4 3.25 86
HOPS-361S 478.99 1.10 70.4 18.8 $\pm$ 3.2 0.23 1.2E+23 4.8E+06 0.66 $\pm$ 0.23 43.1 $\pm$ 3.8 3.50 65
HOPS-384 1477.95 1.98 93.3 51.2 $\pm$ 15.6 0.28 1.4E+23 5.8E+06 0.50 $\pm$ 0.33 30.8 $\pm$ 2.2 2.50 98
HOPS-403 4.14 0.89 22.3 19.0 $\pm$ 3.0 0.77 3.8E+23 1.6E+07 0.80 $\pm$ 0.03 / 2.00 10
HOPS-404 0.95 0.53 20.4 16.0 $\pm$ 0.7 0.53 2.6E+23 1.1E+07 0.37 $\pm$ 0.23 / 1.00 11
Others
HOPS-12W 7.31 0.48 24.8 18.2 $\pm$ 1.4 0.30 1.5E+23 6.1E+06 0.57 $\pm$ 0.18 39.7 $\pm$ 5.0 2.00 23
HOPS-87S / 0.07 / 21.2 $\pm$ 1.8 0.06 2.8E+22 1.2E+06 0.76 $\pm$ 0.18 / 1.00 0
HOPS-88 15.81 0.70 30.0 21.3 $\pm$ 2.1 0.34 1.7E+23 7.1E+06 0.83 $\pm$ 0.25 35.2 $\pm$ 4.3 1.00 35
HOPS-164 0.58 0.24 20.2 15.9 $\pm$ 1.2 0.19 9.7E+22 4.0E+06 0.63 $\pm$ 0.25 / 1.00 7
HOPS-224 2.99 0.36 21.5 13.6 $\pm$ 0.5 0.34 1.7E+23 7.1E+06 0.59 $\pm$ 0.26 / 1.00 6
HOPS-250 6.79 0.20 24.3 14.4 $\pm$ 0.9 0.16 7.8E+22 3.3E+06 0.70 $\pm$ 0.24 / 1.00 5
HOPS-317N 4.76 0.11 22.9 14.7 $\pm$ 0.8 0.09 4.5E+22 1.9E+06 0.37 $\pm$ 0.19 / 1.25 5
HOPS-325 6.20 0.28 23.9 24.6 $\pm$ 5.0 0.22 1.1E+23 4.6E+06 0.44 $\pm$ 0.04 / 1.00 12
HOPS-341 2.07 0.24 21.0 16.0 $\pm$ 0.8 0.23 1.1E+23 4.8E+06 0.19 $\pm$ 0.12 / 1.25 12
HOPS-373E 5.32 0.53 23.2 17.8 $\pm$ 1.5 0.43 2.2E+23 9.0E+06 0.28 $\pm$ 0.18 28.9 $\pm$ 3.6 1.50 33
HOPS-373W 5.32 0.64 23.2 17.8 $\pm$ 1.5 0.52 2.6E+23 1.1E+07 0.50 $\pm$ 0.16 / 1.50 15
HOPS-398 1.01 0.62 20.4 19.8 $\pm$ 2.2 0.55 2.8E+23 1.2E+07 0.50 $\pm$ 0.19 32.0 $\pm$ 5.0 1.00 20
HOPS-399 6.34 1.79 24.0 20.8 $\pm$ 2.3 1.28 6.3E+23 2.6E+07 0.42 $\pm$ 0.15 46.4 $\pm$ 3.2 1.00 103
HOPS-402 0.55 0.51 20.2 14.9 $\pm$ 1.1 0.51 2.5E+23 1.1E+07 0.66 $\pm$ 0.17 / 1.00 10
HOPS-408 0.52 0.31 20.2 13.9 $\pm$ 0.6 0.27 1.3E+23 5.6E+06 / / 1.00 8
OMC1N-4-5-EN / 0.47 / 36.0 $\pm$ 8.7 0.18 9.0E+22 3.8E+06 0.65 $\pm$ 0.26 16.4 $\pm$ 1.4 1.50 66
OMC1N-6-7 / 0.95 / 35.5 $\pm$ 8.8 0.38 1.9E+23 7.8E+06 0.80 $\pm$ 0.58 39.2 $\pm$ 3.8 1.25 53
OMC1N-8-S / 0.16 / 37.0 $\pm$ 10.3 0.06 3.0E+22 1.3E+06 0.48 $\pm$ 0.21 / 1.50 0
Scattering
HH212M 14.00 0.52 29.1 17.2 $\pm$ 0.7 0.32 1.6E+23 6.6E+06 0.39 $\pm$ 0.15 / 1.00 /
HOPS-10 3.33 0.17 21.8 17.6 $\pm$ 0.6 0.12 6.2E+22 2.6E+06 0.49 $\pm$ 0.12 / 1.00 /
HOPS-12E 7.31 0.33 24.8 18.2 $\pm$ 1.4 0.21 1.0E+23 4.3E+06 0.51 $\pm$ 0.23 / 2.50 /
HOPS-50 4.20 0.32 22.4 23.8 $\pm$ 1.9 0.23 1.2E+23 4.8E+06 0.64 $\pm$ 0.29 / 1.00 /
HOPS-53 26.42 0.37 34.1 36.8 $\pm$ 4.0 0.15 7.6E+22 3.2E+06 0.36 $\pm$ 0.15 / 1.00 /
HOPS-60 21.93 0.55 32.6 27.5 $\pm$ 4.7 0.24 1.2E+23 5.0E+06 0.92 $\pm$ 0.30 / 1.00 /
HOPS-81 1.24 0.14 20.5 22.1 $\pm$ 1.6 0.11 5.6E+22 2.4E+06 0.56 $\pm$ 0.20 / 1.00 /
HOPS-84 49.11 0.59 39.8 21.7 $\pm$ 1.7 0.20 1.1E+23 4.2E+06 0.39 $\pm$ 0.23 / 2.00 /
HOPS-153 4.43 0.22 22.5 15.1 $\pm$ 1.1 0.16 7.8E+22 3.3E+06 0.87 $\pm$ 0.27 / 1.00 /
HOPS-203N 20.44 0.42 32.0 16.0 $\pm$ 1.1 0.18 9.0E+22 3.8E+06 0.50 $\pm$ 0.19 / 2.50 /
HOPS-203S 20.44 0.38 32.0 16.0 $\pm$ 1.1 0.16 8.1E+22 3.4E+06 0.43 $\pm$ 0.20 / 2.00 /
HOPS-247 3.09 0.37 21.6 13.4 $\pm$ 0.7 0.34 1.7E+23 7.1E+06 0.74 $\pm$ 0.30 / 1.00 /
HOPS-340 1.85 0.21 20.8 16.0 $\pm$ 0.8 0.20 1.0E+23 4.2E+06 0.36 $\pm$ 0.22 / 1.25 /
HOPS-354 6.57 0.34 24.2 16.2 $\pm$ 0.1 0.18 9.1E+22 3.8E+06 0.54 $\pm$ 0.23 / 1.00 /
HOPS-358 24.96 0.44 33.7 14.7 $\pm$ 1.0 0.22 1.1E+23 4.6E+06 1.34 $\pm$ 0.56 / 1.00 /
HOPS-383 7.83 0.22 25.2 23.4 $\pm$ 3.0 0.14 6.9E+22 2.9E+06 0.70 $\pm$ 0.34 / 1.00 /
Column 2 to 12 present the bolometric luminosity ($L_{\rm{bol}}$), total continuum flux density at 345 GHz ($S_{\nu}$), averaged envelope dust temperature ($\overline{T}_{\rm env}$), cloud dust temperature ($T_{\rm c}$), envelope gas mass ($M_{\rm env}$), gas column density ($N_{\rm gas}$), gas number density ($n_{\rm gas}$), non-thermal velocity dispersion ($\sigma_{\rm nth}$), B-field angle dispersion ($\delta\phi$), fragmentation level ($N_{F}$), and number of Nyquist-sampling B-field segments ($N_{B}$), respectively. “Standard”, “Rotated”, and “Spiral” indicate protostars with different types of B-field structure of standard hourglass, rotated hourglass, and spiral configuration, respectively. “Others” means the B-field pattern is complex or insufficient to identify its morphology. “Scattering” represents protostars with compact polarized emission, which is thought to be produced by self-scattering (Huang et al., 2024).$\dagger$\$$\dagger$\$footnotemark: Typical uncertainties for the different parameters are: 10% for $S_{\nu}$; 20–30% for $\overline{T}_{\rm env}$; 30–40% for $M_{\rm env}$, $N_{\rm gas}$, and $n_{\rm gas}$.
Table 1: Envelope and disk parameters for the BOPS sample\$$\dagger$\$$\dagger$\$footnotemark:
In the case of a dusty cloud heated by internal stellar radiation, the radial temperature profile can be expressed as a power law: $T_{\rm dust}\propto r^{-2/(4+\beta)}$ (Kenyon et al., 1993), where $\beta$ is the dust opacity spectral index. Girart et al. (2009) adopted a value of $\beta=1.5$ when studying low-mass protostars at 1000 au scales. Kwon et al. (2009) suggested that $\beta$ should be around or less than 1 for Class 0 protostars. Tobin et al. (2020) fitted a temperature power-law index of -0.46 for hundreds-au-size disks within the Orion star-forming protostellar system, corresponding to $\beta\sim 0.35$, which is scaled by luminosity. However, this profile shows a slight decline in the outer 100 au regions due to backheating from the surrounding envelope (Tobin et al., 2020), it is subsequently recommended to refine the index to -0.50 (Sheehan et al., 2022). For this study, we use the temperature model proposed by Tobin et al. (2020) for Orion star-forming regions, with $\beta$ set to 1 (corresponding to the index of $\sim-0.40$) for BOPS protostellar envelopes (most of them are in Class 0 stage). Therefore, the dust temperature $T_{\rm dust}$ can be expressed as:
$T_{\rm dust}=T_{0}\biggl{(}\frac{L_{\rm bol}}{L_{\odot}}\biggr{)}^{0.25}\biggl{(}\frac{r}{50~{}{\rm au}}\biggr{)}^{-0.40}$ (2)
where $T_{0}=43$ K is the average temperature for a $\sim$ 1 $L_{\odot}$ protostar at a radius of $\sim$50 au (Whitney et al., 2003; Tobin et al., 2013, 2020), $r$ is the distance from the protostar, and $L_{\rm bol}$ is the bolometric luminosity taken from Furlan et al. (2017). Given the clear detection of C17O emission in our sample, it is inferred that the dust temperature cannot fall below 20 K, as CO would otherwise be frozen onto the dust grains, making it undetectable (e.g., Jørgensen et al., 2015). The Equation 2 may yield very low dust temperatures (below 20 K) at the edge of the selected regions, especially when the bolometric luminosity $L_{\rm bol}$ is very low. In such cases, we use 20 K as the minimum dust temperature for pixels where the calculated temperature falls below this value. Excluding the extremely high temperature in the central pixel, we calculate the average envelope dust temperature $\overline{T}_{\rm env}$ within inner regions with a radius of $R=1200$ au.
Lombardi et al. (2014) used Planck and Herschel observations of Orion to construct and fit SEDs, and further to obtain a map of dust temperature. As shown in Figure 1, comparing the cloud dust temperature $T_{\rm c}$ (Lombardi et al., 2014) with the envelope dust temperature $\overline{T}_{\rm env}$ (this work, derived using our method described above), we find $\overline{T}_{\rm env}$ in most BOPS protostars is larger than $T_{\rm c}$, while 10 protostars have values of $\overline{T}_{\rm env}$ comparable to the dust temperature of the surrounding molecular cloud ($|\overline{T}_{\rm env}-T_{\rm c}|\lesssim 3$, the value of 3 is the half of the mean error for $|\overline{T}_{\rm env}-T_{\rm c}|$, see the notemark of Table LABEL:Tab:paras or Section 4 for error discussions). Given that the scales of molecular clouds in question are on the parsec level, which is much larger than the envelope scales we focus on, it is expected that $\overline{T}_{\rm env}$ should be comparable to or higher than $T_{\rm c}$. However, HOPS-409 shows that the value of $\overline{T}_{\rm env}$ is lower than that of $T_{\rm c}$, probably due to the cloud providing shielding from the interstellar radiation field. We note that 6 protostars lack information of $\overline{T}_{\rm env}$, in this case, $T_{\rm c}$ is employed to estimate the physical properties of the envelope instead.
Figure 1: The difference between derived dust temperature within envelope $\overline{T}_{\rm env}$ and cloud dust temperature $T_{\rm c}$ obtained from Lombardi et al. (2014). Grey dashed line indicates when $\overline{T}_{\rm env}=T_{\rm c}$.
The gas column density $N_{\rm gas}$ is estimated using the relation:
$N_{\rm gas}=\frac{M_{\rm env}}{\pi R^{2}\mu_{\rm gas}m_{\rm H}}$ (3)
where $R=1200$ au is the radius, $M_{\rm env}=\eta M_{\rm dust}$ is the envelope gas mass, $\eta=100$ is the gas-to-dust mass ratio, $\mu_{\rm gas}=2.37$ is the mean molecular weight per free particle (Kauffmann et al., 2008), and $m_{\rm H}$ is the mass of the hydrogen atom. For a spherical region, the average gas number density can be calculated as
$n_{\rm gas}=\frac{3N_{\rm gas}}{4R}$ (4)
Table LABEL:Tab:paras summarizes the derived parameters for each protostar, including information from the literature. Column 2 to 8 of Table LABEL:Tab:paras provide the following data: bolometric luminosity $L_{\rm bol}$, total flux density $S_{\nu}$, mean envelope dust temperature $\overline{T}_{\rm env}$, cloud dust temperature $T_{\rm c}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, and gas number density $n_{\rm gas}$. Figure 2(A–B) illustrate the distributions of bolometric luminosity and envelope dust temperature for 55 protostars (the remaining 6 lack luminosity information), while Figure 2(C–F) present the distributions for parameters of flux intensity, envelope mass, gas column density and number density for 61 targets.
Figure 2: The stacked histogram of physical envelope parameters for protostars with different B-field structure. Panel A to H: bolometric luminosity $L_{\rm bol}$, mean dust envelope temperature $\overline{T}_{\rm env}$, total flux density $S_{\nu}$, envelope gas mass $M_{\rm env}$, gas column density $N_{\rm gas}$, gas number density $n_{\rm gas}$, non-thermal velocity dispersion $\sigma_{\rm nth}$ and angle dispersion of B-field $\delta\phi$, respectively. Dashed lines indicate the mean values of the distribution for protostars with different types of B-field structure.
3.2 Line Width
Assuming that non-thermal motions on envelope scales arise from turbulence, we use the line width of the C17O (3–2) to estimate the turbulence. C17O has hyperfine lines, but they can be ignored because the hyperfine splitting is minimal ($\sim 0.0012$ km s-1) compared to the spectral resolution ($\sim 0.87$ km s-1). To determine the velocity dispersion, we correct the observed velocity dispersion by deconvolving it from the minor contribution from the velocity resolution (e.g., Huang et al., 2023):
$\sigma_{\rm v}^{2}=\sigma_{\rm obs}^{2}-\frac{\Delta v_{\rm r}^{2}}{8\rm ln2}$ (5)
where $\sigma_{\rm v}$ is the velocity dispersion along the line of sight, and $\Delta v_{\rm r}$ is the spectral resolution of the observations $\sim 0.87\ \rm km\ s^{-1}$. The observed velocity dispersion $\sigma_{\rm obs}=\Delta V_{\rm FWHM}/(2\sqrt{2\rm ln2})$, where $\Delta V_{\rm FWHM}$ is estimated using the moment one map of C17O within the same regions as that for the estimation of the physical parameters. In light of the potential impact of the gravity and outflows on observed dispersion, we exculded the innermost regions with a radius of 200 au around the dust peak, and the region where the CO outflow emission is larger than 5-$\sigma$. Then the non-thermal component of the velocity dispersion can be expressed as:
$\sigma_{\rm nth}^{2}=\sigma_{\rm v}^{2}-\sigma_{\rm th}^{2}=\sigma_{\rm v}^{2}-\frac{k_{\rm B}T_{\rm kin}}{m_{\rm obs}}$ (6)
where $\sigma_{\rm nth}$ and $\sigma_{\rm th}$ are the non-thermal and the thermal velocity dispersion, respectively. $k_{\rm B}$ is the Boltzmann constant, and $m_{\rm obs}$ is the mass of the observed molecule (for C17O, $m_{\rm obs}=29\ m_{\rm H}$). $T_{\rm kin}$ is the kinetic temperature of the gas, which we assume to be equal to $\overline{T}_{\rm env}$. The non-thermal velocity dispersions are listed in column 9 of Table LABEL:Tab:paras. Figure 2(G) shows the distribution of $\sigma_{\rm nth}$ for 60 protostars (HOPS-408 has no detection of velocity dispersion within the selected region).
3.3 Angle Dispersion of B-field
The conventional approach for estimating the angle dispersion of B-fields ($\delta\phi$) is by measuring its standard deviation. However, this approach is limited by the 180 ambiguity in polarization angle, which complicates the accurate estimation of the angle dispersion. To address this challenge, further methods have been proposed to quantify $\delta\phi$. Heitsch et al. (2001) proposed a variation in which the restriction for large angles was removed. Later Hildebrand et al. (2009) and Houde et al. (2009) applied a more sophisticated technique which uses an analytical derivation of the dispersion in polarization angles within a turbulent molecular cloud. In a recent study, Cortés et al. (2024) proposed an alternative approach to derive the angle dispersion by assuming that the local turbulence perturbing the field follows a Gaussian distribution, and they suggested that 16 Nyquist-sampling segments of B-field are required for the Gaussian distribution under the assumption of 5° for the margin of error and 10° for the mean standard deviation of polarization angles. In this paper, we follow the typical method to estimate the angle dispersion $\delta\phi$ by calculating the standard deviation of the distribution of B-field segments. In order to reduce the ambiguity associated with the 180° orientation of the B-field, the distribution of the position angles is shifted when the angles are centred near 0° or 180°. The standard deviation of the shifted distribution is regarded as $\delta\phi$. To avoid the contamination of self-scattering (Huang et al., 2024), the polarization segments within the innermost 400 au region ($\sim 1\arcsec$, corresponding to the beam size) have been masked. We also mask the polarization data for regions beyond a radius of 1200 au, consistent with the scales used to derive the physical parameters in Section 3.1. Protostellar envelopes with less than 16 Nyquist-sampling segments of B-field within 1200 au regions have large uncertainties ($>5^{\circ}$) according to Cortés et al. (2024), therefore we only estimate the angle dispersion of B-field for 25 protostars with sufficient statastics ($>$ 16 Nyquist-sampling segments), which is listed in column 10 of Table LABEL:Tab:paras. Among these 25 protostellar envelopes, 6 exhibit B-field morphology that is standard-hourglass, 8 are rotated-hourglass, 4 are spiral, and 7 are complex or lack polarized emission to identify their B-field morphology. Figure 2(H) shows the distribution of $\delta\phi$ for each type of B-field morphology.
3.4 Fragmentation Level and Disk Properties
[…]
Figure 4: The stacked histogram of disk parameters for protostars with different B-field structure. Upper row: disk mass $M_{\rm disk}$ (panel A), disk radius $R_{\rm disk}$ (B). Bottom row: disk inclination angle $\theta_{\rm Inc}$ (panel C and D). Dashed lines in A, B and C panels indicate the mean values of the distribution for different B-field structures. In panel D, “Face-on”, “Inclined”, and “Edge-on” indicate the inclination angle within a range of 0–30, 30–60, and 60–90, respectively. The disk mass, disk radius, and disk inclination angle are obtained from Tobin et al. (2020).
Tobin et al. (2020) identified protostellar disks in the Orion molecular clouds using both VLA and ALMA observations, with further refinements presented in their subsequent study of Tobin et al. (2022). All objects in our sample are included within the VANDAM survey, which enables us to estimate the multiplicity of these systems. Palau et al. (2013) defined the fragmentation level as the number of millimeter sources inside a region with a radius of 0.05 pc for massive dense cores, whereas Busquet et al. (2016) used a different radius of 0.15 pc. We follow the above methods, but define a different radius of 1200 au, which is the scale for estimating the physical parameters above. Then the number of fragments (including small-scale fragmentation, that was only resolved in the survey with higher resolution of $\sim$40–400 au-scale companions) identified from the VANDAM and the BOPS projects within the innermost envelope region is indicative of the fragmentation level $N_{\rm F}$. It is possible that the outer ($>1200$ au) regions may influence the fragmentation of the envelopes, given that the typical scale for the star-forming envelopes is up to several thousand au (e.g., Myers et al., 2000). We thus define larger radii of 1800 au and 2400 au for estimating extra contribution of $N_{\rm F}$, but with less weighting values of 0.5 and 0.25, respectively. For example, in Figure 3, which shows the 0.87 mm dust emission map of an example HOPS-78, a protostar (black asterisk) is located in the innermost region with a radius of 1200 au (black dashed circle), then $N_{\rm F}$ in this region is counted as 1. While there are 2 (blue asterisks) and 1 (red asterisk) protostars in annular regions with inner and outer radii of 1200 and 1800 au (blue dashed circle), and 1800 and 2400 au (red dashed circle), then the $N_{\rm F}$ for each region is counted as 1 and 0.25, respectively. Note that these protostars are identified from continuum maps of the BOPS and VANDAM survey. The total fragmentation level $N_{\rm F}$ for HOPS-78 is therefore assigned a value of 2.25. Column 11 of Table LABEL:Tab:paras lists the fragmentation level for 61 targets.
The mass, radius, size, and inclination angle of the disks listed in column 5–8 of Table LABEL:Tab:parameters in the Appendix A are taken from Tobin et al. (2020). Figure 4 shows the distribution of 43 protostars. The rest 18 protostars lack information of the disk, or the disk dust emission is potentially contaminated by the envelope. The upper panels show the distribution of the disk mass and radius, while the bottom panel presents the distribution of disk inclination angle, which is estimated by $\theta_{\rm inc}=\arccos(\Theta_{\rm min}/\Theta_{\rm maj})$, where $\Theta_{\rm maj}$ and $\Theta_{\rm min}$ are the deconvolved major and minor axes of the disk. Here the face-on and edge-on configurations are defined with an inclination angle of $\theta_{\rm Inc}$ = 0–30, and $\theta_{\rm Inc}$ = 60–90, respectively. The trend of the distribution for disk properties, as well as other parameters, will be discussed in the following.
4 Discussion
In Section 3, we have derived physical parameters of envelope gas mass, gas column density, gas number density for entire sample of BOPS (61 protostars) based on the available bolometric luminosity (Furlan et al., 2017) and dust temperature (Lombardi et al., 2014), and non-thermal velocity dispersion for 60 protostars within the inner region of 6″ (corresponding to a scale of 2400 au, with a radius of 1200 au). Angle dispersion of B-field is also estimated within the same regions for 25 BOPS targets with sufficient Nyquist-sampling statistics. Combining with the VANDAM survey, the fragmentation level has been identified for 61 BOPS targets. In addition, disk properties have been presented for 43 BOPS targets from Tobin et al. (2020), without any contamination of envelope emission and other components of disk emission.
The uncertainties of the derived physical parameters are due to unknowns and errors in factors such as dust opacities, dust temperatures, and observed flux. The uncertainty in physical parameters resulting from dust opacity, an internally consistent systematic error, does not significantly impede the identification of correlations between physical parameters, as these are calculated in a uniform manner, which translates the distributions up or down. Whereas the dust temperature and the flux calibration are considered random. For the temperature, based on the model that we used, we consider typical uncertainties of 20–30%, and we expect flux calibration to be accurate within 10%. All combined, we derive uncertainties of about 30–40% for the masses and densities.
The distributions of the these parameters are shown in Figure 2. Combining both the distribution and the mean values (dashed lines) for each type of B-field morphologies, we find that protostars with high bolometric luminosity ($>$ 50 L) and/or envelope dust temperature ($>$ 35 K) tend to exhibit spiral or rotated-hourglass B-field morphology (Figure 2(A–B)). In Figure 2(C–F), envelopes with compact polarized emission (self-scattering in blue) tend to have lower envelope masses, gas column densities, and gas number densities. Whereas protostars exhibiting a standard-hourglass B-field pattern tend to exhibit relatively larger values in these parameters. However, these standard-hourglass protostars have relatively low non-thermal velocity dispersion, which is potentially indicative of the magnetic braking effect, whereby the strong field exerts a braking influence on the rotational rate (as shown in panel G of Figure 2). In Figure 2(H), we find standard-hourglass-B-field protostars have relatively low angle dispersion, in contrast to rotated-hourglass-, spiral-, and complex-field sources, which display a relatively larger angle dispersion.
In Figure 4, which shows the distribution of disk properties, we find from the first row (panel A and B) that protostars exhibiting standard hourglass B-field structure have smaller disk size than those exhibiting other field configurations. In panel C, the mean values for protostars with different type of B-field structure are ranging from 25 to 50. Note that the finite resolution and vertical thickness of the disks (Sheehan et al., 2022; Lin et al., 2023; van’t Hoff et al., 2023) will bias this simple inclination estimate against 90, which would influence the location of the mean values. In panel D, in both types of sources dominated by self-scattering and exhibiting rotated-hourglass morphologies, with the mean polarization angle preferentially aligned parallel to the outflow direction, the disk inclination angle is found to be either very inclined or edge-on. However, in standard-hourglass sources, with the mean polarization angle oriented perpendicularly to the outflow, the disk inclination angle is typically found to be face-on.
Below we discuss the potential correlation between these parameters and disk properties. To statistically examine the relations between two pairs of parameters, we performed both the Pearson linear correlation coefficient ($r_{p}$) and the Spearman’s rank correlation coefficient ($r_{s}$). The values of $r_{p}$ and $r_{s}$ are each bounded from -1 to 1, for negative to positive slope correlations, respectively. In each case, the closer the absolute value is to 1, the stronger the correlation. The p–value is used for the purpose of testing the null hypothesis that the two variables are uncorrelated. In general, statistically significant correlations are identified by p–value $<$ 0.05. Here we analyze the following figures, and based on these details, we will motivate a more detailed analysis on fragmentation and the possible effect of magnetic braking on disk formation.
4.1 Physical Envelope Properties
Figure 5: Left panels: Distributions between the bolometric luminosity $L_{\rm bol}$ and the envelope gas mass $M_{\rm env}$ (panel A), the gas column density $N_{\rm gas}$ (C), the gas volumn density $n_{\rm gas}$ (E), and the non-thermal velocity dispersion $\sigma_{\rm nth}$ (G). Right panels: Distribution between the mean dust temperature $\overline{T}_{\rm env}$ and the envelope gas mass $M_{\rm env}$ (B), the gas column density $N_{\rm gas}$ (D), the gas volume density $n_{\rm gas}$ (F), and the non-thermal velocity dispersion $\sigma_{\rm nth}$ (H).
The dust temperature is expected to be correlated with the bolometric luminosity if the dust is primarily heated by the stellar radiation (see Section 3.1). During the early stages of star formation (Class 0 and early Class I), the protostar is deeply embedded within a dense envelope. In this phase, the luminosity is predominantly due to the accretion of material from the envelope through the disk onto the protostar, which releases gravitational energy. High-density regions are often associated with intense accretion activity as the protostar accumulates material. The inflow of material from the envelope onto the protostar leads to an increase in $L_{\rm bol}$. Additionally, a more massive envelope absorbs and re-radiates more energy and it is more efficient at cooling due to its higher optical depth. 5(A–F) shows the distribution between these physical parameters for a sample of 55 envelopes (the remaining 6 targets have no bolometric luminosity information). However, our results reveal no significant correlations between $L_{\rm bol}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, and between $\overline{T}_{\rm env}$ and $M_{\rm env}$, $N_{\rm gas}$, $n_{\rm gas}$, as their correlation coefficients are close to 0, with p–values much higher than 0.05. The lack of direct correlations is likely because protostellar envelope masses are influenced by both the final stellar mass of the protostar and the evolutionary stage, with envelope masses decreasing rapidly over time (Fischer et al., 2017). Luminosity, on the other hand, is influenced by episodic accretion, which may vary independently of the envelope mass (e.g., Zakri et al., 2022; Narang et al., 2024).
Non-thermal motions, often driven by turbulence, are of significant importance in star-forming regions and affect the overall dynamics and evolution of the protostellar system. Therefore, $\sigma_{\rm nth}$ serves as an indicator of turbulence level, which enhance the accretion process by increasing the rate at which material is fed onto the protostar. Protostars with higher value of $L_{\rm bol}$ typically drive powerful outflows and winds, injecting energy into the surrounding medium and thereby increasing the non-thermal velocity dispersion. In high turbulent regions, the non-thermal velocity dispersion provide support against gravitational collapse. Once, however, collapse is triggered, the turbulent motions channel material onto the forming protostar, influencing its luminosity. Theoretical models also suggest that higher levels of turbulence ($\sigma_{\rm nth}$) lead to more efficient accretion processes, leading to increased protostellar luminosities (e.g., Krumholz et al., 2005; Padoan et al., 2011; Federrath, 2015). Additionally, in regions with high $\sigma_{\rm nth}$, which are characterised by intense star formation activity (e.g., jets and winds), one might expect that higher dust temperatures may result from the combined effects of shock heating and enhanced radiative feedback from very young protostars. To test these senarios, we plot the distribution between $\sigma_{\rm nth}$ and $L_{\rm bol}$, and between $\sigma_{\rm nth}$ and $T_{\rm env}$ for 54 protostars (5 protostars lack luminosity information, while one protostar of HOPS-408 has no detection of velocity dispersion), as shown in Figure 5(G–H). A slight positive correlation is observed between $\sigma_{\rm nth}$ and $L_{\rm bol}$ from $r_{s}$ and p-value, as well as between $\overline{T}_{\rm env}$ and $L_{\rm bol}$ from $r_{p}$ and $r_{s}$. However, $r_{p}\sim 0$ in panel G and the p-value $>$ 0.05 in panel H suggest insignificant correlations between these parameters. The interplay between localized feedback, external turbulence, B-fields, and accretion variability makes it difficult to isolate a direct relationship, especially when dealing with deeply embedded young protostars in Orion, a highly turbulent environment. In addition, the limited spectral resolution would also affect these results. More focused studies on individual cores or higher-resolution observations might reveal more localized correlations that are otherwise masked by these factors.
We also note that these results will be influenced by the assumption of a typical temperature of 20 K. However, we find that 7 protostars exhibit relatively high values $\overline{T}_{\rm env}$ ($>$ 40 K). Of these, 2 display a rotated hourglass field shape (HOPS-288 and HOPS-370), 5 of the protostars are spiral (HOPS-124, HOPS-182, HOPS-361N, HOPS-361S, and HOPS-384).
4.2 Fragmentation
Figure 6(A–C) show the relation between $N_{\rm F}$ and $M_{\rm env}$ for 61 protostars, between $N_{\rm F}$ and $\sigma_{\rm nth}$ for 60 protostars (HOPS-408 has no C17O detection, so its $\sigma_{\rm nth}$ cannot be resolved), and between $N_{\rm F}$ and $\delta\phi$ for 18 protostars out of 25 with enough polarzation statistics (the other 7 have a complex large-scale B-field morphology), respectively. It is important to note that the values of $N_{\rm F}$ are dominated by multiples within the inner 2400 au region, a scale that is consistent with the one used to estimate the physical parameters as discussed in Section 3. Combined with the statistical test methods, we discuss the potential influence of specific factors on envelope fragmentation at the 2400 au scale in the following.
Optical and near infrared studies suggest that the multiplicity rate drops with decreasing mass (e.g., Chini et al., 2012, 2013), we thus investigated the correlation between $N_{\rm F}$ and $M_{\rm env}$. Based on the correlation coefficients shown in Figure 6(A), there is no clear correlation between $N_{\rm F}$ and $M_{\rm env}$. This result may suggest that the correlation observed in the optical/infrared is not yet established at the very early stages of star and cluster formation. It is possible, therefore, that the final number of members in a multiple system is determined after multiple star formation episodes, as suggested by both theoretical (e.g., Myers, 2011) and observational(e.g., Bik et al., 2012) studies. Furthermore, the process of fragmentation and the formation of multiple objects are typically governed by a complex interplay of factors beyond just the envelope mass. These interactions may lead to significant variability in outcomes, masking any simple correlation between envelope mass and multiplicity. Previous observational studies of massive dense cores also suggest that there is no correlation between mass and multiplicity within a radius of 0.05 pc of the dust peak (e.g., Palau et al., 2013).
In recent decades, several studies on the turbulence scenario (e.g., Padoan et al., 2001; Vázquez-Semadeni et al., 2007) have demonstrated that turbulence is capable of generating density structures through the action of supersonic shocks, which compress and fragment the gas efficiently, particularly in the presence of significant turbulence. Subsequent studies have indicated that different types of turbulence, i.e. compressive and solenoidal, exert divergent effects on fragmentation (e.g., Federrath et al., 2008). Compressive turbulence is conducive to fragmentation, whereas solenoidal turbulence has the effect on slowing down fragmentation (e.g., Federrath et al., 2010; Girichidis et al., 2011). As shown in Figure 6(B), the fragmentation level $N_{\rm F}$ is uncorrelated with $\sigma_{\rm nth}$. However, it is important to note that $\sigma_{\rm nth}$ may include contributions from rotational and infalling motions on envelope scales, which cannot be distinguished with the limited spectral resolution of C17O. Therefore, deeper studies with sufficient spectral resolution should be carried out to reach more robust conclusions.
On the other hand, the fragmentation level in massive dense cores could be explained by a B-field/turbulence balance (Palau et al., 2013) or changes in the mass-to-flux ratio (Busquet et al., 2016; Añez-López et al., 2020; Palau et al., 2021) on scales of several thousands au, suggesting a potential correlation between B-fields and fragmentation on envelope scales. One of the most longstanding methods for estimating the B-field strength is the Davis-Chandrasekhar-Fermi (DCF, Davis, 1999; Chandrasekhar et al., 1953) method, which relies on the principle that turbulent motions should visibly affect the distribution of the B-field in the plane of the sky. The basic form of the DCF formula is $B_{\rm{pos}}=\sqrt{4\pi\rho}\sigma_{\rm nth}/\delta\phi$, where $\rho$ is the gas density, and $\delta\phi$ is the measured angle dispersion of the B-field (as discussed in Section 3.3). Since $\sigma_{\rm nth}$ is barely resolved for most sources in our sample, which significantly affect the estimation of B-field strength. However, the above equation indicates that the field strength is inversely proportional to the dispersion in the polarization position angles, which provides a potential way to indirectly study the correlation between field strength and other physical parameters on envelope scales. In order to statistically test this correlation, we plot the distribution between the angle dispersion of the B-field $\delta\phi$ and the fragmentation level $N_{\rm F}$, as shown in Figure 6(C). Combing correlation coefficients, we find that $N_{\rm F}$ is positively correlated with $\delta\phi$. This suggests that higher angle dispersions in the B-field is associated with higher multiplicities, and thus envelope fragmentation tends to be suppressed by the B-field. Previous studies have indicated that strong B-fields suppress fragmentation, as they constitute an additional form of support against gravitational contraction and also via the magnetic effect (inward pinching and gathering of the field lines dragged during collapse makes the magnetic strength larger at inner radius, and acts against rotation), which reduces the amount of angular momentum in the inner part of cores (e.g., Galli et al., 2006; Mellon et al., 2008; Hennebelle et al., 2008; Price et al., 2009; Myers et al., 2014; Cunningham et al., 2018; Kwon et al., 2019; Yen et al., 2021).
In addition, multiplicity may also decrease with the evolution from Class 0 to Class I (Tobin et al., 2020, 2022). However, since the BOPS sample predominantly consists of objects at the Class 0 stage, the sample size in our study is insufficient to differentiate evolutionary factors over time.
Figure 6: Panel A–C: Distributions between the fragmentation level $N_{\rm F}$ and the envelope mass $M_{\rm env}$ (panel A), the non-thermal velocity dispersion $\sigma_{\rm nth}$ (panel B), and the angle dispersion of B-field $\delta\phi$ (panel C). Panel D–F: Distributions between the disk radius $R_{\rm disk}$ and the alignment of the B-field and outflow axes $|\Delta\theta|$ (panel D), and between the disk mass $M_{\rm disk}$ and $|\Delta\theta|$ (panel E and F). In panel F, disk masses with their disk radii less than 100 au are derived from Sheehan et al. (2022).
4.3 B-field and Disk Formation
Observations indicate that cloud cores in star-forming environments are magnetized (e.g., Crutcher, 1999). Theoretical models suggest that B-field tension forces deflect infalling gas towards the equatorial plane, resulting in a dense, flattened pseudo-disk (e.g., Allen et al., 2003). During the collapse phase, the field lines are drawn inward into a pinched geometry (e.g., Girart et al., 2006, 2009; Stephens et al., 2013; Kwon et al., 2019), which strengthens the B-field and carries the angular momentum outward. This process is known as magnetic braking, which reduce angular momentum by one order of magnitude, potentially suppressing the formation of a centrifugally supported disk (e.g., Galli et al., 2006). Figure 6(D) presents the distribution between the disk size $R_{\rm disk}$ and the B-field angle dispersion ($\delta\phi$, which is inversely proportional to the B-field strength) for 11 protostars with good detection in the disk radius. We find that protostellar envelopes with a standard hourglass B-field structure tend to have a smaller disk radius and angle dispersion of the B-field than those with a rotated hourglass field structure, suggesting that magnetic braking is more effective for the standard hourglass field shape. Some studies propose that an initial misalignment between the B-field and the rotation axis would lead to weaker magnetic braking of the collapsing core, enabling earlier disk formation (e.g., Joos et al., 2012; Li et al., 2013; Hirano et al., 2019; Machida et al., 2020).
Huang et al. (2024) have revealed the dust polarized continuum emission and outflow signatures in protostars, finding that 40% of the protostars exhibit a mean B-field direction that is approximately perpendicular to the outflow on envelope scales. However, the relative orientation in the remaining sample appears to be random. We combine the disk radius taken from Tobin et al. (2020) and envelope misalignment between the B-field and outflow for BOPS protostars, as shown in Figure 6(E). Protostars with complex and unresolved polarized emission, as well as those with disks significantly affected by envelope emission and other component of disk emission (see Table LABEL:Tab:parameters in Appendix A and figures in Appendix B) are excluded in our analysis. This exclusion is due to the substantial uncertainty in the averaged B-field direction or disk radius/mass for these sources. Then the number of sources used in Figure 6(E) and 6(F) is 16. In panel E, protostars with standard hourglass, rotated hourglass, and spiral B-field structures exhibit a strong positive correlation between disk radius $R_{\rm disk}$ and the misalignment angle between magnetic and outflow axes $|\Delta\theta|$.
Sheehan et al. (2022) found that the disk masses derived from Tobin et al. (2020) are overestimated compared to radiative transfer model, with the discrepancy becoming more pronounced for lower-mass disks. To address this, they proposed a log-linear relation to reconcile the disk masses derived from Tobin et al. (2020) with those obtained through radiative transfer modelling, expressed as:
$\log(M_{\rm disk}^{\rm Sheehan})=\frac{\log(M_{\rm disk})-(0.83\pm 0.05)}{(0.67\pm 0.03)}$ (7)
Figure 6(F) shows the distribution between the re-estimated disk mass and $|\Delta\theta|$. A similar discernible trend can be found from this panel that $M_{\rm disk}$ is positively correlated with $|\Delta\theta|$, but the correlation is weaker compared to the $R_{\rm disk}$-$|\Delta\theta|$ relation. If the bottom-left point in Figure 6(E–F) is excluded, the updated correlation coefficients are $r_{p}\sim 0.74$, $r_{s}\sim 0.75$, p–value $<0.01$ for panel E, and $r_{p}\sim 0.62$, $r_{s}\sim 0.59$, p–value $\sim 0.02$ for panel F, respectively. These results indicate that the correlations between the disk radius/mass and the misalignment angle remain consistent, even when the bottom-left point is excluded, suggesting that the correlation is less important affected by this point. The p-values, which are $<$ 0.05, suggest that these correlations are statistically significant. In addition, it is important to note from the distribution of both panel E and panel F that protostars exhibiting standard hourglass B field tend to have lower disk size, compared to those exhibiting spiral and rotated hourglass field structure.
Huang et al. (2024) applied both the intensity-weighted method (which gives more weight to the inner protostellar envelope) and the error-weighted method (which provides an averaged weight across the selected envelope) to estimate the mean direction of the (B-field). The mean direction distributions of B-field obtained from both weighting methods were found to be comparable throughout the sample, suggesting that the measured B-field is minimally influenced by gravitational collapse and outflows. In this case, the measured B-field is likely representative of the original pre-stellar core B-field. Furthermore, the outflow direction aligns with the angular momentum of the original core. Our findings suggest that a greater misalignment between these two directions lead to weaker magnetic braking of the collapsing core and result in a more prominent disk, which is consistent with the misalignment-induced disk formation model (e.g., Joos et al., 2012).
5 Conclusions
We have calculated and analyzed the physical parameters for 61 protostelar envelopes within the BOPS sample. The combination of the envelope properties derived from the BOPS (Huang et al., 2024) and the disk properties derived from the VANDAM survey (Tobin et al., 2020) allows us to link the turbulence and B-field in the envelope with the resulting fragmentation and disk properties. The main results are as follows:
  1. 1.
    Protostellar envelopes exhibiting standard-hourglass field shape show higher values of envelope mass, gas column density, and gas number density than those with other field configurations, lower non-thermal velocity dispersion have also be observed in this type. In contrast, envelopes with a spiral B-field structure tend to have higher non-thermal velocity dispersion, indicating a more turbulent environment.
  2. 2.
    The angle dispersion of the B-field, which is inversely proportional to the B-field strength, is positively correlated with the fragmentation level, suggesting that the pinched field suppresses the fragmentation, as it provides additional support against gravitational collapse, and/or brakes the rotation efficiently. However, the lack of correlations between the fragmentation level and the envelope mass, and between the fragmentation level and the non-thermal velocity dispersion, may be due to the complex interplay of various factors influencing fragmentation beyond just the envelope mass, or to the limited spectral resolution.
  3. 3.
    Protostars with standard hourglass B-field structure have smaller disk size and angle dispersion of the magnetic field than those with other field configurations, specially rotated hourglasses, but also spirals and others. This suggests that in the case of the standard hourglass field structure, the magnetic field strength is stronger, ultimately leading to a smaller disk through the more effective magnetic braking, in line with theory predictions.
  4. 4.
    A more significant misalignment between the magnetic and outflow axes is associated with larger disk radii and disk masses. This suggests that increased misalignment results in weaker magnetic braking during core collapse, which aligns with the misalignment-induced disk formation model.
  5. 5.
    The dust temperature and bolometric luminosity are not correlated with envelope masses and densities, probably because the cloud with low bolometric luminosity provides shielding from the interstellar radiation field. In addition, no clear correlation can be found between non-thermal velocity dispersion and luminosity on envelope scales, contrast with the simulation models predicted. Deeper studies on individual cores or higher-resolution observations might reveal more localized correlations that are otherwise masked by the larger-scale dynamics.
Insignificant correlations found for some pairs of parameters, especially in the analysis of the non-thermal velocity dispersion, may be due to the limited sensitivity and resolution of the observations. It is important to note that the non-thermal velocity dispersion may include contributions from disk rotation, envelope rotation, and/or infalling motions, further observations with higher sensitivity and resolution should be carried out to distinguish them and further refine our conclusions about the non-thermal velocity dispersion or the turbulence.
B.H. and J.M.G. acknowledge support by the grant PID2020-117710GB-I00 (MCI-AEI-FEDER, UE). B.H. also acknowledges financial support from the China Scholarship Council (CSC) under grant No. 202006660008. This work is also partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M. M.F.L. acknowledges support from the European Research Executive Agency HORIZON-MSCA-2021-SE-01 Research and Innovation programme under the Marie Skłodowska-Curie grant agreement number 101086388 (LACEGAL). M.F.L. also acknowledges the warmth and hospitality of the ICE-UB group of star formation. L.W.L. acknowledges support by NSF AST-1910364 and NSF AST-2307844. Z.Y.L. is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. W.K. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00342488). The authors acknowledge Anaëlle Maury and Chat Hull for helpful discussions, Jacob Labonte for early analysis of the BOPS data. This paper makes use of the following ALMA data: ADS/JAO.ALMA#2019.1.00086. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.
Appendix A Table
Table 2: Some parameters for the BOPS sample
Name R.A. Decl. $D$ $M_{\rm disk}$ $R_{\rm disk}$ $\langle\Theta_{\rm max}\times\Theta_{\rm min}\rangle_{\rm Decon.}$ $\theta_{\rm Inc}$ $\Delta\theta$
(h:m:s) (d:m:s) (pc) (M) (au) (${}^{\prime\prime}\times^{\prime\prime}$) () ()
Standard
HOPS-11 05:35:13:43 -05:57:57.91 388.3 135.3 $\pm$ 2.4 41.9 $\pm$ 10.0 0.13 $\times$ 0.12 22.6 14.0 $\pm$ 9.7
HOPS-87Naafootnotemark: 05:35:23.42 -05:01:30.57 392.7 / / / / /
HOPS-359 05:47:24.84  00:20:59.39 429.4 89.8 $\pm$ 2.0 62.0 $\pm$ 10.0 0.17 $\times$ 0.15 28.1 15.6 $\pm$ 7.2
HOPS-395 05:39:17.09 -07:24:24.59 397.2 16.9 $\pm$ 0.9 7.1 $\pm$ 10.0 0.02 $\times$ 0.02 0 0.1 $\pm$ 9.0
HOPS-400 05:42:45.26 -01:16:13.89 415.4 106.6 $\pm$ 4.7 48.7 $\pm$ 10.0 0.14 $\times$ 0.13 21.3 14.3 $\pm$ 6.6
HOPS-407 05:46:28.25  00:19:27.95 419.1 83.2 $\pm$ 2.4 64.1 $\pm$ 10.0 0.18 $\times$ 0.14 38.9 16.5 $\pm$ 7.8
OMC1N-8-Nccfootnotemark: 05:35:18.20 -05:20:48.57 392.8 / / / / /
Rotated
HH270IRS 05:51:22.72  02:56:05:00 405.7 161.3 $\pm$ 3.4 92.3 $\pm$ 10.0 0.27 $\times$ 0.14 58.8 68.3 $\pm$ 9.6
HOPS-78 05:35:25.96 -05:05:43.38 392.8 187.9 $\pm$ 2.1 147.1 $\pm$ 10.0 0.44 $\times$ 0.14 71.4 86.0 $\pm$ 6.6
HOPS-168bbfootnotemark: 05:36:18.94 -06:45:23.58 383.3 / / / / /
HOPS-169 05:36:36.16 -06:38:54.41 384.0 118.5 $\pm$ 2.5 53.8 $\pm$ 10.0 0.17 $\times$ 0.13 40.1 51.3 $\pm$ 7.2
HOPS-288bbfootnotemark: 05:39:56.00 -07:30:27.62 405.5 / / / / /
HOPS-310 05:42:27.68 -01:20:01.35 414.3 263.8 $\pm$ 5.3 101.3 $\pm$ 10.0 0.29 $\times$ 0.15 58.9 89.9 $\pm$ 6.4
HOPS-317Saafootnotemark: 05:46:08.38 -00:10:43.59 427.1 / / / / /
HOPS-370 05:35:27.64 -05:09:34.40 392.8 114.6 $\pm$ 2.2 112.8 $\pm$ 10.0 0.34 $\times$ 0.11 71.1 86.8 $\pm$ 5.9
HOPS-401 05:46:07.73 -00:12:21.31 426.9 185.9 $\pm$ 4.2 74.7 $\pm$ 10.0 0.21 $\times$ 0.16 40.4 78.8 $\pm$ 7.6
HOPS-409 05:35:21.37 -05:13:17.88 392.8 110.7 $\pm$ 1.1 112.1 $\pm$ 10.0 0.34 $\times$ 0.12 69.3 73.3 $\pm$ 6.1
OMC1N-4-5-ESccfootnotemark: 05:35:15.97 -05:20:14.23 392.8 / / / / /
Spiral
HOPS-96 05:35:29.72 -04:58:48.63 392.7 135.0 $\pm$ 2.3 41.4 $\pm$ 10.0 0.12 $\times$ 0.11 23.6 52.4$\pm$ 6.2
HOPS-124ababfootnotemark: 05:39:19.91 -07:26:11.22 398.0 / / / / /
HOPS-182 05:36:18.79 -06:22:10.24 385.1 35.7 $\pm$ 1.3 36.3 $\pm$ 10.0 0.11 $\times$ 0.06 56.9 32.7 $\pm$ 6.3
HOPS-303 05:42:02.65 -02:07:45.94 410.0 206.8 $\pm$ 4.4 82.9 $\pm$ 10.0 0.24 $\times$ 0.19 37.7 53.9 $\pm$ 12.2
HOPS-361Nbbfootnotemark: 05:47:04.63  00:21:47.82 430.4 / / / / /
HOPS-361S 05:47:04.78  00:21:42.79 430.4 138.6 $\pm$ 1.4 96.1 $\pm$ 10.0 0.26 $\times$ 0.15 54.8 40.2 $\pm$ 11.0
HOPS-384bbfootnotemark: 05:41:44.14 -01:54:46.00 409.5 / / / / /
HOPS-403aafootnotemark: 05:46:27.91 -00:00:52.15 428.2 / / / / /
HOPS-404aafootnotemark: 05:48:07.72  00:33:51.81 430.1 / / / / /
Others
HOPS-12W 05:35:08.63 -05:55:54.65 388.6 20.1 $\pm$ 0.4 20.5 $\pm$ 10.0 0.06 $\times$ 0.05 33.6 /
HOPS-87Sccfootnotemark: 05:35:23.67 -05:01:40.27 392.7 / / / / /
HOPS-88 05:35:22.47 -05:01:14.33 392.7 70.5 $\pm$ 1.2 34.0 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-164 05:37:00.43 -06:37:10.91 385.0 133.5 $\pm$ 2.8 70.6 $\pm$ 10.0 0.22 $\times$ 0.14 50.5 /
HOPS-224 05:41:32.07 -08:40:09.82 440.3 236.6 $\pm$ 5.0 107.7 $\pm$ 10.0 0.29 $\times$ 0.18 51.6 /
HOPS-250 05:40:48.85 -08:06:57.11 428.5 133.9 $\pm$ 2.9 117.2 $\pm$ 10.0 0.32 $\times$ 0.19 53.6 /
HOPS-317N 05:46:08.60 -00:10:38.49 427.1 51.6 $\pm$ 1.8 34.5 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-325 05:46:39.20  00:01:12.30 428.5 58.6 $\pm$ 1.1 34.9 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-341 05:47:00.92  00:26:21.50 430.9 28.3 $\pm$ 1.1 54.5 $\pm$ 10.0 0.15 $\times$ 0.09 53.1 /
HOPS-373E 05:46:31.10 -00:02:33.05 428.1 79.2 $\pm$ 1.8 69.8 $\pm$ 10.0 0.19 $\times$ 0.17 26.5 /
HOPS-373W 05:46:30.91 -00:02:35.16 428.1 70.3 $\pm$ 0.8 36.0 $\pm$ 10.0 0.10 $\times$ 0.09 25.8 /
HOPS-398 05:41:29.41 -02:21:16.39 408.0 271.8 $\pm$ 6.0 76.9 $\pm$ 10.0 0.22 $\times$ 0.19 30.3 /
HOPS-399 05:41:24.93 -02:18:06.66 407.9 196.7 $\pm$ 4.3 66.2 $\pm$ 10.0 0.19 $\times$ 0.15 37.9 /
HOPS-402aafootnotemark: 05:46:10.03 -00:12:16.92 426.9 / / / / /
HOPS-408 05:39:30.90 -07:23:59.75 398.9 124.5 $\pm$ 3.6 93.5 $\pm$ 10.0 0.28 $\times$ 0.24 31.0 /
OMC1N-4-5-ENccfootnotemark: 05:35:16.05 -05:20:05.75 392.8 / / / / /
OMC1N-6-7ccfootnotemark: 05:35:15.70 -05:20:39.31 392.8 / / / / /
OMC1N-8-Sccfootnotemark: 05:35:18.00 -05:20:55.77 392.8 / / / / /
Scattering
HH212M 05:43:51.41 -01:02:53.20 413.2 102.9 $\pm$ 2.3 78.6 $\pm$ 10.0 0.22 $\times$ 0.10 63.0 /
HOPS-10 05:35:09.05 -05:58:26.89 388.2 52.5 $\pm$ 1.0 46.8 $\pm$ 10.0 0.14 $\times$ 0.07 60.0 /
HOPS-12Ebbfootnotemark: 05:35:08.95 -05:55:54.99 388.6 / / / / /
HOPS-50 05:34:40.92 -05:31:44.74 391.5 172.8 $\pm$ 3.3 82.8 $\pm$ 10.0 0.25 $\times$ 0.13 58.7 /
HOPS-53 05:33:57.40 -05:23:30.00 390.5 33.4 $\pm$ 0.5 21.6 $\pm$ 10.0 0.07 $\times$ 0.05 44.4 /
HOPS-60 05:35:23.29 -05:12:03.45 392.8 152.7 $\pm$ 2.7 62.7 $\pm$ 10.0 0.19 $\times$ 0.11 54.6 /
HOPS-81 05:35:28.01 -05:04:57.36 392.8 66.4 $\pm$ 1.2 32.4 $\pm$ 10.0 0.10 $\times$ 0.06 53.1 /
HOPS-84 05:35:26.56 -05:03:55.15 392.8 159.2 $\pm$ 3.0 87.4 $\pm$ 10.0 0.26 $\times$ 0.12 62.5 /
HOPS-153 05:37:57.03 -07:06:56.27 387.9 103.4 $\pm$ 2.1 129.1 $\pm$ 10.0 0.39 $\times$ 0.10 75.1 /
HOPS-203Nbbfootnotemark: 05:36:22.87 -06:46:06.63 383.5 / / / / /
HOPS-203S 05:36:22.90 -06:46:09.54 383.5 33.1 $\pm$ 0.4 30.9 $\pm$ 10.0 0.10 $\times$ 0.06 53.1 /
HOPS-247 05:41:26.19 -07:56:51.90 430.9 227.0 $\pm$ 4.6 85.3 $\pm$ 10.0 0.23 $\times$ 0.16 45.9 /
HOPS-340 05:47:01.31  00:26:23.05 430.9 59.9 $\pm$ 1.0 41.7 $\pm$ 10.0 0.11 $\times$ 0.07 50.5 /
HOPS-354 05:54:24.26  01:44:19.87 355.4 30.3 $\pm$ 0.6 24.4 $\pm$ 10.0 0.08 $\times$ 0.07 29.0 /
HOPS-358 05:46:07.26 -00:13:30.25 426.8 110.7 $\pm$ 1.0 136.3 $\pm$ 10.0 0.38 $\times$ 0.10 74.7 /
HOPS-383 05:35:29.79 -04:59:50.38 392.8 96.0 $\pm$ 1.9 61.7 $\pm$ 10.0 0.18 $\times$ 0.12 48.2 /
Column 2 to 9 present the R.A., Decl., distance ($D$), disk mass ($M_{\rm disk}$), disk radius ($R_{\rm disk}$), deconvolved disk size ($\langle r_{\rm max}\rangle\times\langle r_{\rm min}\rangle$), disk inclined angle ($\theta_{\rm Inc}=\arccos(\langle\Theta_{\rm min}\rangle/\langle\Theta_{\rm max}\rangle)$), and the difference between outflow and B-field direction ($\Delta\theta$), respectively. The source distance, disk mass, disk radius, disk size and their uncertainty are obtained from Tobin et al. (2020). The difference between outflow and B-field is obtained from Huang et al. (2024), which is the alignment between the mean outflow direction and the uncertainty-weighted mean B-field direction.aafootnotemark: The disk emission is contaminated by the envelope dust emission.bbfootnotemark: The disk emission is contaminated by other disk emission.ccfootnotemark: No information on disk properties presented.
Table 2: Envelope and disk parameters for the BOPS sample\$$\dagger$\$$\dagger$\$footnotemark:
Appendix B Figures
B.1 Self-scattering dominated protostars
Left panels: the velocity field in color scale (obtained from the C17O (3–2) line) overlaid with the 870 $\mu$m dust polarization segments and Stokes I contours. Blue/red contours indicate the blue-/red-shifted outflow, with counter levels set at 5 times the outflow rms × (1, 2, 4, 8, 16, 32). The black contour levels for the Stokes I image are 10 times the rms × (1, 2, 4, 8, 16, 32, 64, 128, 256, 512). For the velocity field, regions with an S/N less than 4 have been flagged. The purple segments represent the polarization. The magenta dashed circle corresponds to scales with a radius of 1200 au. Right panels: 9 mm dust maps obtained from Tobin et al. (2020).
B.2 Aligned dust grain dominated protostars
Left panels: the velocity field in color scale (obtained from the C17O (3–2) line) overlaid with the 870 $\mu$m dust polarization segments and Stokes I contours. Blue/red contours indicate the blue-/red-shifted outflow, with counter levels set at 5 times the outflow rms × (1, 2, 4, 8, 16, 32). The black contour levels for the Stokes I image are 10 times the rms × (1, 2, 4, 8, 16, 32, 64, 128, 256, 512). For the velocity field, regions with an S/N less than 4 have been flagged. The white segments represent the magnetic fields. The dashed magenta circle corresponds to scales with a radius of 1200 au. The purple curves in categories of standard hourglass, rotated hourglass, and spiral indicate the B-field patterns. Right panels: 9 mm dust maps obtained from the VANDAM survey (Tobin et al., 2020).
Standard hourglass [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
Rotated hourglass [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]
Spiral [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
Others
[Uncaptioned image]              [Uncaptioned image]              [Uncaptioned image]
Others (continue) [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]